### v1

# Diffusion Dataset Condensation: Training Your Diffusion Model Faster with Less Data

Abstract
Diffusion models have achieved remarkable success in various generative tasks, but training them remains highly resource-intensive, often with millions of images and GPU days of computation required. From a data-centric perspective addressing the limitation, we study diffusion dataset condensation as a new challenging problem setting that aims at constructing a “synthetic” sub-dataset with significantly fewer samples than the original dataset for training high-quality diffusion models significantly faster. To the best of our knowledge, we are the first to formally study the dataset condensation task for diffusion models, while conventional dataset condensation focused on training discriminative models. For this new challenge, we further propose a novel Diffusion Dataset Condensation (D2C) framework, that consists of two phases: Select and Attach. The Select phase identifies a compact and diverse subset via a diffusion difficulty score and interval sampling, upon which the Attach phase enhances conditional signals and information of the selected subset by attaching rich semantic and visual representations. Extensive experiments across dataset sizes, model architectures, and resolutions demonstrate that our D2C can train diffusion models significantly faster with dramatically fewer data while retaining high visual quality. Notably, for the SiT-XL/2 architecture, our D2C achieves a 100$\times$ acceleration, reaching a FID of 4.3 in just 40k steps using only 0.8% of the training data.
footnotetext: Equal contribution.
   Correspondence to zekexie@hkust-gz.edu.cn.
Figure 1: Our D2C framework significantly accelerates diffusion model training with limited data. (a) Overview of our D2C pipeline, which consists of a Select phase that filters a compact and diverse subset via diffusion difficulty score and interval sampling, and an Attach phase that enriches samples with semantic and visual information. (b) D2C achieves over 100$\times$ faster convergence compared to REPA and over 233$\times$ faster than vanilla SiT-XL/2, reaching a FID of 4.3 at just 40k steps. (c) Under a strict 4% data budget (0.05M), our method achieves a FID of 2.7 at 180k iterations, demonstrating its strong training efficiency and rapid convergence.
1 Introduction
Generative models, such as score-based [1, 2, 3, 4] and flow-based [5, 6, 7] approaches, have achieved remarkable success in various generative tasks, producing high-quality and diverse data across domains [8, 9, 10]. However, these approaches are notoriously data and compute intensive to train, often requiring millions of samples and hundreds of thousands of iterations to capture complex high-dimensional distributions [11, 12, 13]. The resulting cost presents a significant barrier to broader application and iteration within the AIGC community, making efficient training increasingly important across both academic and industrial settings [14, 15]. Recent efforts have improved diffusion training efficiency through various strategies, such as architectural redesigns [13, 16, 11], attention optimization [17], reweighting strategies [18, 19, 20, 8], and representation learning [12]. In parallel, data-centric approaches such as Patch Diffusion [21] and advanced augmentation techniques [8] aim to better exploit the potential of existing data. Despite these advancements, the exploration of building a relatively complete “synthetic” subset through dataset condensation [22] to accelerate diffusion model training with minimal information loss remains underexplored.
Dataset condensation [22, 23, 24, 25] aims to construct a “synthetic” sub-dataset with significantly fewer samples than the original dataset, such that a model trained from scratch on this subset achieves comparable performance to one trained on the full dataset. Unlike data pruning or selection [26], which passively select existing samples, condensation actively optimizes synthetic data, offering greater potential for aggressive data reduction and training efficiency [27]. However, all existing methods are designed for discriminative tasks. Compared to discriminative tasks, generative tasks are much more complex and demand higher dataset quality [28]. Applying popular methods (e.g., SRe2[24]) that have substantiated effective for discriminative tasks to diffusion models presents significant challenges, such as the failure to produce diverse, high-quality outputs with structural and semantic fidelity, leading to degraded results and unstable convergence (see Sec. 4).
We raise a key question: “Can we train diffusion models dramatically faster with significantly less data, while retaining high generation quality?” The answer is affirmative. Answering this question holds significant relevance for the further development of visually generative intelligence and is therefore extremely worthwhile to explore. In this paper, we mainly made the following contributions.
First, to the best of our knowledge, we are the first to formally study the dataset condensation task for diffusion models, a new challenging problem setting that aims at constructing a “synthetic” sub-dataset with significantly fewer samples than the original dataset for training high-quality diffusion models significantly faster. More specifically, our explorations with the diffusion model provide the first insights into the challenges and potential solutions for applying dataset condensation to vision generation tasks. We note that while conventional dataset condensation made great progress and sometimes use diffusion models to construct a subset, this line of research only focused on training discriminative models instead of generative models.
Second, we propose a a novel two-stage Diffusion Dataset Condensation (D2C) framework, consisting of Select and Attach, illustrated in Fig. 1 (a). Our framework addresses the challenges of dataset condensation for diffusion models by decomposing the problem into two key aspects: the Select stage identifies an informative, compact, and learnable subset by ranking samples using the diffusion difficulty score derived from a pre-trained diffusion model; the Attach stage enriches each selected sample by adding semantic and visual representations, further enhancing the training efficiency while preserving performance.
Third, extensive experiments demonstrate that the proposed D2C can train diffusion models significantly faster with dramatically fewer data while retaining high visual quality, substantiating the effectiveness and scalability. Specifically, D2C significantly outperforms random sampling and several popular dataset selection and distillation algorithms (e.g., SRe2[24] and K-Center) across data compression ratios of 0.8%, 4%, and 8%, at resolutions of 256$\times$256 and 512$\times$512, and with both SiT [13] and DiT [11] architectures. In particular, D2C achieves a FID of 4.3 in merely 40k training steps (w/o classifier-free guidance (CFG) [4]) using SiT-XL/2 [13], demonstrating a 100$\times$ acceleration over REPA [12] and a 233$\times$ speed-up compared to vanilla SiT. Furthermore, it further improves to a FID of 2.7 using only 50k synthesized images with CFG (refer to Fig. 1 (c)).
2 Preliminaries and Related Work
In this section, we briefly review diffusion models as well as dataset condensation.
Diffusion Models. We briefly introduce the standard latent-space noise injection formulation that underlies many diffusion models [6, 11, 9]. These models define a forward process that gradually perturbs input data with Gaussian noise, and train a neural network to approximate the reverse denoising process for sampling (see Appendix B for more details).
Let the original variable be denoted as $\mathbf{x}_{0}\sim q_{0}(\mathbf{x})$. The forward diffusion process is typically implemented as a Markov chain with Gaussian noise injections:
$q_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}),$ (1)
where $\alpha_{t},\ \sigma_{t}\in\mathbb{R}^{+}$ are differentiable functions of $t$ with bounded derivatives. The choice for $\alpha_{t}$ and $\sigma_{t}$ is referred to as the noise schedule of a diffusion model. A commonly used reparameterization in score-based diffusion models [3] expresses the noisy data $\mathbf{x}_{t}$ at time step $t$ as $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\epsilon$, where $\ \epsilon\sim\mathcal{N}(0,\mathbf{I})$. In this case, to recover “clean” data from noise, a neural network $\epsilon_{\theta}(\cdot,\cdot,\cdot)$ is trained to predict the added noise $\epsilon$, thereby approximating the reverse process. This framework adopt the conventional noise prediction strategy for training, where the training objective is to minimize the mean squared error between the predicted and the ground true noise:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\|^{2}_{2}\right],$ (2)
Here, $\mathbf{c}$ is a conditional input, such as class labels or text embeddings. In some cases, the prediction target is replaced with a linear combination of noise and data (e.g. $v$-prediction), which corresponds to flow matching [5, 6, 7]:
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-(\epsilon-\mathbf{x}_{0})\|^{2}_{2}\right].$ (3)
Here, $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ denotes the learned time-dependent velocity field.
Data-centric Efficient Training. Various model-side strategies have been proposed to accelerate diffusion model, including architectural enhancements [17, 11, 13], sampling refinements [18, 19, 20, 8], and representation-level techniques that leverage pretrained vision features [12, 29]. However, relatively little has been explored from a data-centirc perspective. In data-centric model training, given an original dataset $\mathcal{D}=(\hat{\mathbf{X}},\hat{\mathbf{Y}})=\{(\hat{\mathbf{x}}_{i},\hat{y}_{i})\}_{i=1}^{|\mathcal{D}|}$, where each $\hat{y}_{i}$ is the label corresponding to sample $\hat{\mathbf{x}}_{i}$, dataset compression aims to reduce the size of training data while preserving model performance. Two primary strategies have been extensively studied in this context: dataset pruning and dataset condensation.
1) Dataset Pruning. Dataset pruning selects an information-enrichment subset from the original dataset, i.e., $\mathcal{D}^{\text{core}}\subset\mathcal{D}$ with $|\mathcal{D}^{\text{core}}|\ll|\mathcal{D}|$, and directly minimizes the training loss over the subset:
$\min_{\theta}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}^{\text{core}}}\left[\ell(\phi_{\theta_{\mathcal{D}^{\text{core}}}}(\mathbf{x}),y)\right],$ (4)
where $\ell(\cdot,\cdot)$ denotes the empirical training loss, and $\phi_{\theta_{\mathcal{D}^{\text{core}}}}$ is the model parameterized by $\theta_{\mathcal{D}^{\text{core}}}$. Classical data pruning methods like random sampling, K-Center [30], and Herding [31] can be used with diffusion models, but they offer minimal performance improvements.
2) Dataset Condensation. In contrast, dataset condensation aims to synthesize a small, compact, and diverse synthetic dataset $\mathcal{D}^{\mathcal{S}}=(\mathbf{X},\mathbf{Y})=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{|\mathcal{D}^{\mathcal{S}}|}$ to replace the original dataset $\mathcal{D}$. The synthetic dataset $\mathcal{D}^{\mathcal{S}}$ is generated by a condensation algorithm $\mathcal{C}$ such that $\mathcal{D}^{\mathcal{S}}\in\mathcal{C}(\mathcal{D})$, with $|\mathcal{D}^{\mathcal{S}}|\ll|\mathcal{D}|$. Each $y_{j}$ corresponds to the synthetic label for the sample $\mathbf{x}_{j}$.
The key motivation for dataset condensation is to create $\mathcal{D}^{\mathcal{S}}$ such that models trained on it can achieve performance within an acceptable deviation $\eta$ compared to models trained on $\mathcal{D}$. This can be formally expressed as:
$\sup\left\{\left|\ell(\phi_{\theta_{\mathcal{D}}}(\hat{\mathbf{x}}),\hat{y})-\ell(\phi_{\theta_{\mathcal{D}}^{\mathcal{S}}}(\hat{\mathbf{x}}),\hat{y})\right|\right\}_{(\hat{\mathbf{x}},\hat{y})\sim\mathcal{D}}\leq\eta,$ (5)
where $\theta_{\mathcal{D}}$ is the parameter set of the neural network $\phi$ optimized on $\mathcal{D}$: $\theta_{\mathcal{D}}=\arg\min_{\theta}\mathbb{E}_{(\hat{\mathbf{x}},\hat{y})\sim\mathcal{D}}\left[\ell(\phi_{\theta}(\hat{\mathbf{x}}),\hat{y})\right].$ A similar definition applies to $\theta_{\mathcal{D}}^{\mathcal{S}}$, which is optimized on the synthetic dataset $\mathcal{D}^{\mathcal{S}}$. Existing methods, such as RDED, MTT, SRe2L, and G-VBSM [27, 24, 25, 32, 33], are primarily designed for discriminative tasks. When applied to diffusion models, these methods generate synthetic images that deviate from the original data distribution, leading to detrimental effects on model training. Visualization of these synthesized images can be found at Appendix I.
Figure 2: Overview of Diffusion Dataset Condensation (D2C). D2C employs a two-stage process: Select and Attach. The Select stage identifies a compact and diverse subset by intervaling sampling using the diffusion difficulty score derived from a pre-trained diffusion model. The Attach stage further enriches each selected sample by adding semantic information and visual information.
3 Diffusion Dataset Condensation
To enable data-centric efficient training of diffusion models under limited resources, we propose Diffusion Dataset Condensation (D2C), the first unified framework that systematically condenses training data for diffusion models. As illustrated in Fig. 2, this process produces a condensed dataset suitable for efficient diffusion model training. D2C consists of two stages: Select (Sec. 3.1), which identifies a compact set of diverse and learnable real images using diffusion difficulty score and interval sampling techniques; and Attach (Sec. 3.2), which augments each selected image with semantic and visual information to improve generation performance. Finally, we present the novel training paradigm utilizing the condensed dataset generated by D2C (Sec. 3.3).
3.1 Select: Difficulty-Aware Selection
In this work, we focus on class-to-image (C2I) synthesis. Given a class-conditioned dataset $\mathcal{D}=\bigcup_{y=1}^{C}\mathcal{D}_{y}$, where $\mathcal{D}_{y}=\{x_{i}\}_{i=1}^{|\mathcal{D}_{y}|}$ denotes all samples of class $y$, we aim to select a compact subset for efficient diffusion training. To achieve this, we propose diffusion difficulty score to quantify the denoising difficulty of each sample, followed by our designed interval sampling to ensure diversity within the selected subset.
Diffusion Difficulty Score. The arrangement of samples from easy to hard is crucial for revealing underlying data patterns and facilitating difficulty-aware selection. Recent work [34] demonstrates that diffusion models inherently encode semantic-related class-conditional probability $p_{\theta}(\mathbf{c}|\mathbf{x})$ through the variational lower bound (i.e., diffusion loss Eq. 2) of $\log p_{\theta}(\mathbf{x}|\mathbf{c})$ [3, 1]. This conditional probability can be formulated as $p_{\theta}(\mathbf{c}|\mathbf{x})=\frac{p_{\theta}(\mathbf{x}|\mathbf{c})p(\mathbf{c})}{\sum_{\mathbf{\hat{c}}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})p(\mathbf{\hat{c}})}$. Intuitively, a larger $p_{\theta}(\mathbf{c}|\mathbf{x})$ indicates that sample $\mathbf{x}$ can be more confidently identified as belonging to class $\mathbf{c}$, thus suggesting lower learning difficulty. Given the significant computational overhead of the full Bayesian formulation and our focus on estimating sample difficulty, we ignore the denominator part $\sum_{\mathbf{\hat{c}}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})p(\mathbf{\hat{c}})$ of the calculation. Since the category $y\sim U\{0,\cdots,C\}$ ($C$ denotes the class number) is obtained by uniform sampling, and assuming $\sup\{|\mathbb{E}_{\mathbf{\hat{c}}}[p_{\theta}(\mathbf{x_{1}}|\mathbf{\hat{c}})]-|\mathbb{E}_{\mathbf{\hat{c}}}[p_{\theta}(\mathbf{x_{2}}|\mathbf{\hat{c}})]|\}_{\mathbf{x_{1}},\mathbf{x}_{2}\sim\mathcal{D^{X}}}\leq\eta$ ($\mathcal{D^{X}}$ denotes the all images in the dataset), we define the diffusion difficulty score based on the class-conditional probability $p_{\theta}(\mathbf{x}|\mathbf{c})\propto\frac{p_{\theta}(\mathbf{x}|\mathbf{c})}{\sum_{\mathbf{c}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})}=p_{\theta}(\mathbf{c}|\mathbf{x})$ ($\sum_{\mathbf{c}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})$ can be viewed as a constant when sorting by sample difficulty):
$s_{\text{diff}}(\mathbf{x})=-p_{\theta}(\mathbf{x}|\mathbf{c})=-\log\left(\exp\left(-\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\left\|\epsilon-\epsilon_{\theta}\left(\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon,\,t,\,\mathbf{c}\right)\right\|^{2}_{2}\right]\right)\right),$ (6)
The higher the score $s_{\text{diff}}(\mathbf{x})$, the easier it is, and the lower the score $s_{\text{diff}}(\mathbf{x})$, the more difficult it is. Note that in this paper’s figures and tables, we directly use the diffusion loss (i.e., $\propto-p_{\theta}(\mathbf{c}|\mathbf{x})$) for ease of understanding.
By computing $s_{\text{diff}}(x)$ for all training samples, we construct a ranked dataset. As shown in Fig. 3 (Left), these scores exhibit a skewed unimodal distribution. Selecting the easiest samples (Min) yields a subset dominated by clean, background-simple images with high learnability but limited diversity. In contrast, selecting only the highest-score samples (Max) results in cluttered, noisy, and ambiguous images that are difficult to optimize. Meanwhile, many samples lie in the middle range, offering moderate learnability but richer contextual information.
Interval Sampling. To balance diversity and learnability, we propose an interval sampling strategy. Specifically, we sort its images $\mathcal{D}_{y}$ within each class $y$ in ascending order of $s_{\text{diff}}(x)$ and select samples at a fixed interval $k$:
$\mathcal{D}_{\text{IS}}=\bigcup_{y=1}^{C}\left\{x^{(i)}\in\mathcal{D}_{y}\;\middle|\;i\in\{0,k,2k,\dots\}\right\},$ (7)
where $\mathcal{D}_{\text{IS}}$ denotes the selected subset constructed by interval sampling, $k$ is the fixed sampling interval, and $x^{(i)}$ is the $i$-th sample in the sorted list (e.g., $x^{(0)}$ corresponds to the sample with the lowest diffusion difficulty score). Interval sampling with a larger interval $k$ promotes diversity in the sampled data while potentially hindering learnability. As shown in Fig. 3 (Left), this trade-off arises from a shift in the sample distribution: a larger $k$ leads to a reduction in the number of easy samples and a corresponding increase in the representation of standard and difficult samples.
Figure 3: Left: Distribution of diffusion difficulty scores under different interval values $k$. Smaller intervals (e.g., 1, 2) favor low-loss samples, while larger intervals (e.g., 64, 128) result in a distribution closer to random sampling, thus approximating the original data distribution. Moderate intervals (e.g., 16) provide balanced coverage across difficulty levels. Right: Representative samples selected by three strategies: Min (lowest score), Max (highest score), and Interval (our proposed strategy). Interval sampling achieves a balance between structural clarity and contextual richness.
Following interval sampling yields notable findings. First, interval sampling efficiently mitigates the inclusion of both overly simplistic and excessively challenging samples, yielding subsets that exhibit both learnability and diversity. Fig. 3 (right) displays representative examples selected by our strategy, showcasing structural clarity coupled with contextual richness. Second, the FID of diffusion models trained on the subset $\mathcal{D}_{\text{IS}}$ generated with increasing $k$ initially decreases and subsequently increases. For example, as shown in Fig. 8, our results further corroborate that $k=16$ strikes a favorable balance between sample diversity and generation quality when using a 50K data budget.
3.2 Attach: Semantic and Visual Information Enhancement
To further enhance the information richness of the selected data, we introduce the Attach phase that augments each instance with semantic and visual information. While the Select phase generates a compact and informative subset, the achievable performance ceiling based solely on this phase remains limited. Consequently, we inject more comprehensive semantic and visual representations into the selected subset to further bolster our method’s generalization capability.
Dual Conditional Embedding (DC-Embedding). Existing C2I synthesis methods [11, 13] commonly rely on class embeddings trained from scratch, which often fail to effectively capture inherent semantic information (see Appendix F.1). We enrich the class embedding by incorporating text representations derived from a pre-trained text encoder (e.g., T5-encoder [35]). For each class $c\in\{1,\dots,C\}$, a descriptive prompt $P(c)$ (e.g., “a photo of a cat”) is encoded by a pre-trained text encoder $f_{\text{text}}$, yielding its corresponding text embedding $t_{c}$ and text mask $t_{\textrm{mask}}$:
$\begin{split}t_{c},t_{\textrm{mask}}&=f_{\text{text}}(P(c)),\\ \end{split}$ (8)
The resulting text embedding and text mask are stored on disk as attached text information alongside the subset $\mathcal{D}_{\textrm{IS}}$ generated in the preceding phase, ready for import during formal training. During the formal training, as illustrated in Fig. 5, the text embedding $t_{c}$ and the text mask $t_{\textrm{mask}}$ undergo a 1D convolution and are fused with a learnable class embedding $e_{c}$ using a residual MLP:
[…]
Figure 4: Visualization of samples from Random (i.e., random sampling) vs. D2C. D2C demonstrates a significant performance advantage over Random across all training iterations, under both a strict 0.8% (10K) and a 4% (50K) data budget, leading to a substantial acceleration in training.
Figure 5: Overview of DC-Embedding.
This resulting vector $y_{\text{text}}$ then serves as a semantic conditioning token for the conditional diffusion model. Compared to using simple class embeddings alone, this formulation offers richer semantic information while retaining the learnability of class embeddings.
Visual Information Injection. While semantic information aids in distinguishing inter-class structure, it often fails to capture the intra-class variability essential for high-fidelity generation. To address this, we integrate instance-specific visual representations into the attached information. For each image $x\in\mathbb{R}^{3\times H\times W}$, a pre-trained vision encoder $f_{\text{vis}}$ (e.g., DINOv2 [36]) extracts patch-level semantic representations:
$y_{\text{vis}}=f_{\text{vis}}(x)\in\mathbb{R}^{N\times d_{\text{text}}}$ (10)
[…]
3.3 D2C Training Process
Here, we detail the training process of the diffusion model using our condensed dataset, which comprises a compact subset selected during the Select phase and subsequently enriched with semantic and visual information during the Attach phase. Our goal is to fully leverage the information existed within our condensed dataset to accelerate training without compromising performance.
We employ a conditional diffusion model $\mathcal{D}_{\theta}$ and, as an example, utilize the optimization objective of score-based diffusion models: predicting the added noise $\epsilon$ from the perturbed latent input $\mathbf{x}_{t}$ at time step $t$, conditioned on the text information $y_{\textrm{text}}$ and the class label $y$. The new denoising loss is defined as $\mathcal{L}^{\text{new}}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,y,y_{\textrm{text}})\|^{2}_{2}\right]$, where the specific injected forms of $y$ and $y_{\text{text}}$ can be found in Sec. 3.2. Then, to maximize the utilization of visual information, we adopt the same formulation as REPA [12], which involves aligning the encoder’s output (i.e., the decoder’s input) within the diffusion model with the visual representation $y_{\text{vis}}=\{v_{i}\}_{i=1}^{h}$. Concretely, from a designated intermediate layer of the diffusion backbone, we obtain token features $\{h_{i}\in\mathbb{R}^{d}\}_{i=1}^{h}$. A projection head $\phi$ maps these tokens from $\mathbb{R}^{d}$ to $\mathbb{R}^{d_{\text{text}}}$, and we compute a semantic alignment loss:
$\mathcal{L}_{\text{proj}}=-\frac{1}{h}\sum_{i=1}^{h}\left\langle\frac{\phi(h_{i})}{\|\phi(h_{i})\|},\frac{v_{i}}{\|v_{i}\|}\right\rangle.$ (11)
This loss encourages the model to align its encoder’s output with visual representations extracted from the visual encoder, promoting localized realism and spatial consistency [36] in generation.
Overall Training Objective. The final training loss combines the denoising objective and the semantic alignment term (with the balance weight $\lambda$ is set to 0.5 by default):
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{diff}}+\lambda\mathbb{E}_{\mathbf{x},\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1],y,y_{\text{text}},y_{\text{vis}}}\left[\mathcal{L}_{\text{proj}}\right].$ (12)
This unified training strategy enables D2C to effectively learn from limited yet enhanced data, offering a practical solution for efficient diffusion training under resource-constrained settings.
Table 1: Comparison of gFID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet 256$\times$256. We use CFG=1.5 for evaluation. D2C surpasses other methods at all settings.
Data Budget Iter. DiT-L/2 SiT-L/2
Random K-Center Herding D2C Random K-Center Herding D2C
0.8% (10K) 100k 35.86 50.77 40.75 4.20 4.35 14.77 22.96 3.98
0.8% (10K) 300k 4.19 13.5 22.35 4.13 4.33 13.58 22.55 3.98
4.0% (50K) 100k 36.78 69.86 32.38 14.81 31.13 61.66 29.11 11.21
4.0% (50K) 300k 11.55 38.54 22.44 5.99 14.18 39.69 22.44 5.66
8.0% (100K) 100k 41.02 71.31 36.37 22.55 36.64 66.96 32.3 15.01
8.0% (100K) 300k 11.49 37.35 15.23 6.49 12.56 39.08 16.17 5.65
Table 2: Comparison with a strict data budget 0.8% (10K) on ImageNet 512$\times$512. We use CFG=1.5 for evaluation. D2C surpasses random sampling at all settings. D2C’s performance on SiT-L/2 with 300k iterations can be found in Appendix J.
[…]
DiT-L/2 Random 100k 24.8 11.9 74.3 0.65 0.42
DiT-L/2 D2C (Ours) 100k 14.8 6.9 109.2 0.63 0.52
DiT-L/2 Random 300k 17.1 12.8 130.6 0.64 0.41
DiT-L/2 D2C (Ours) 300k 5.8 15.1 318.9 0.77 0.29
SiT-L/2 Random 100k 13.3 22.8 197.1 0.69 0.68
SiT-L/2 D2C (Ours) 100k 9.1 14.3 261.7 0.72 0.34
4 Experiments
We validate the performance of D2C and analyze the contributions of its components through extensive experiments. In particular, we aim to answer the following questions:
1) Can D2C improve training speed and reduce data usage of diffusion models? (Table 1, Figs. 14)
2) Does D2C generalize well across backbones, data scales, and resolutions? (Tables 12, Fig. 4)
3) How do D2C’s components and hyperparameter choices affect its overall effectiveness? (Figs. 368)
Experiment settings. We conduct experiments on the ImageNet-1K dataset [37], using subsets of 10K, 50K, and 100K images, corresponding to 0.8%, 4%, and 8% of the full dataset, respectively. All images are center-cropped and resized to 256$\times$256 and 512$\times$512 resolutions using the ADM [16] preprocessing pipeline. Furthermore, we use [$\cdot$]-L/2 and [$\cdot$]-XL/2 architectures in both DiT [11] and SiT [13] backbones, following the standard settings outlined in Ma et al. [13]. More details on implementation and training can be found in Appendix C.
Figure 6: Left: Ablation studies of the interval value $k$ in interval sampling at different data budgets. With a small interval value, training can experience significant acceleration during the early iterations. However, due to the limited diversity of the selected dataset, performance tends to be overtaken by random sampling in later iterations. As the interval value increases, the final gFID-10K scores initially improve and then decline, suggesting an optimal interval value exists. Notably, this optimal value is 96 for a 10K data budget and 16 for a 50K data budget, and the ratio 96/16 closely approximates the ratio of the data budgets, 50K/10K. Right:. Ablation study of DC-Embedding under a 10K data budget. “Only Class” stands for the baseline approach of injecting class embeddings. This figure demonstrates that both the text embedding branch and the original class embedding branch (see Fig. 5) are critical for the effective training of diffusion models.
Table 3: D2C vs. SRe2[24] on ImageNet 256$\times$256 with a strict data budget 0.8% (10K).
[…]
DiT-L/2 SRe2L 104.2 20.2 14.1 0.20 0.30
DiT-L/2 D2C (Ours) 4.2 11.0 283.6 0.72 0.24
SiT-L/2 SRe2L 82.3 19.8 18.1 0.27 0.19
SiT-L/2 D2C (Ours) 3.9 10.7 289.7 0.73 0.25
Evaluation and baselines. We train models from scratch on the collected subset and evaluate them using gFID [38], sFID, Inception Score [39], Precision (Prec.), and Recall (Rec.), adhering to standard evaluation protocols [16, 11, 13]. We compare our method against REPA [12], and various data condensation and selection baselines, including SRe2[24], Herding, K-Center, and random sampling, using SiT and DiT architectures [13, 11]. Further details regarding evaluation metrics and baseline methods can be found in Appendix G and H.
4.1 Main Result
Training Performance and Speed. We evaluate D2C on SiT-XL/2 using 10K and 50K data budgets, comparing its performance against REPA and a vanilla SiT model trained on the full ImageNet dataset (a 1.28M data budget), as well as random selection with 10K and 50K data budgets. As illustrated in Fig. 1 (b), our method achieves a gFID-50K of 4.23 at only 40K iterations with 10K training data and without CFG. In contrast, REPA requires 4 million steps and the vanilla SiT model needs over 7 million steps to reach comparable performance, representing an acceleration of over 100$\times$ and 233$\times$, respectively. Under a 4% data budget (50K) with CFG set to 1.5, our method achieves an FID of 2.78 at 180K steps, further demonstrating significant data and compute efficiency (Fig. 1 (c)). Moreover, Fig. 4 presents a visual comparison between random selection and our D2C at 10K and 50K data sizes, where images are generated from the same noise and class label. Our method demonstrates superior visual quality compared to the baseline and generates higher-quality images, even during the early iterations of training.
Comparison on ImageNet 256$\times$256. We compare D2C with random sampling, Herding [40], K-Center [30], and SRe2[24] under various data budgets and backbones. As illustrated in Table 1, D2C consistently achieves the lowest FID across all settings. For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a gFID-50K of 4.20 on DiT-L/2 and 3.98 on SiT. These results demonstrate the superiority of our approach over existing methods. Notably, SRe2L, which performs well in classification task, fails on this generative task (see Table 3) due to its focus on category-discriminative features. Similarly, geometry-based methods like Herding and K-Center, along with random sampling, prove inadequate for achieving efficient and high-performing training of diffusion models.
Comparison on ImageNet 512$\times$512. As shown in Table 2, D2C achieves a gFID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet 512$\times$512 settings. On SiT-L/2, similar improvements are observed. These results demonstrate that D2C generalizes well to higher resolutions.
4.2 Ablation Study
Figure 7: Ablation sutdies of interval sampling and DC-Embedding. “Mean=N” denotes the selection of a subset of samples where the average diffusion loss is closest to the value N.
Ablation on Select Phase. We investigate the impact of the interval value $k$ in the Select phase, as shown in Fig. 6 (Left) and Fig. 7. Using a small value accelerates early training by prioritizing min-loss samples, which are simpler and easier to learn. However, the limited diversity of such samples leads to degraded performance in later stages, eventually being overtaken by settings with moderate interval values. In contrast, large intervals or random selection introduce excessive max-loss or uncurated samples, destabilizing training (Fig. 3). As $k$ increases, we observe that gFID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability. Empirically, the best results are achieved with an interval of 96 for the 10K budget and 16 for 50K, approximately following the ratio of data budgets (50K/10K).
[…]
DiT-L/2 37.07
DiT-L/2 9.01
DiT-L/2 7.62
Table 4: Ablation on semantic and visual information in the Attach phase.
Ablation on Attach Phase. We conduct ablation studies on the Attach phase from two perspectives. First, as shown in Fig. 6 (Right), we examine the effectiveness of dual conditional embedding under a 10K data budget. Using only class or text embedding results in degraded performance compared to their combination, indicating the necessity of both components to capture category semantics and text embeddings. Notably, text embedding alone performs better than class embedding, suggesting that textual descriptions provide richer semantic information for generative tasks. Second, as shown in Table 4, we analyze the impact of the semantic and visual information injection modules. The gFID-10K increases to 37.07 when both branches are removed and to 9.01 when only the visual injection module is excluded. Notably, the combination of both modules yields the best FID of 7.62. These results underscore the crucial role of injecting both semantic and visual information in accelerating diffusion model training.
5 Conclusion
In this paper, we introduce D2C, the first dataset condensation franework that can significantly accelerate diffusion model training for the generative task. Our pipeline comprises two key phases: Select and Attach. In the Select phase, we leverage diffusion difficulty score and interval sampling to obtain a subset that is both compact and diverse. Subsequently, in the Attach phase, we augment this subset with critical semantic and visual information, leading to a dramatic improvement in the training acceleration ratio and enhanced performance robustness. While there are still some limitations (refer to Appendix A), our D2C as the pioneering work in this novel approach still achieved impressively 100233×100$\sim$ 233$\times$100 ∼ 233 × training acceleration over the baselines. We believe that our work will offer novel inspirations and motivate more research on this promising field in the future.
Broader Impact
We introduce D2C, a method engineered to decrease training overhead while minimizing performance degradation. The dataset employed for training D2C was ImageNet-1K, which adheres to ethical guidelines. Moreover, we prioritize the responsible and ethical deployment of this technology, aiming to maximize its societal benefits while proactively addressing and mitigating any potential risks.
Appendix
Appendix A Limitation
This work primarily focuses on accelerating and improving the training process of diffusion models. While our approach demonstrates promising results in the early stage, the final performance upon convergence does not yet fully match baseline benchmarks. It is important to unleash the potential of our algorithm, such as fine-tuning the vision and text encoders to better align noise latent features with the integrated visual information, thereby reducing existing discrepancies. Additionally, although the current study centers on image generation tasks, extending our method to other domains, such as 3D generation and video synthesis, represents a promising direction for future exploration.
Appendix B Additional Descriptions of Diffusion Models
This section presents the preliminaries of the diffusion model. For simplicity, we introduce the standard denoising diffusion probabilistic model (DDPM), where the forward process gradually corrupts the input with noise and the reverse process restores the target data distribution. We also summarize the architectural details in B.2.
B.1 Denoising Diffusion Probabilistic Model (DDPM)
The DDPM framework models data generation via a discrete-time Markov chain that progressively adds Gaussian noise to a data sample $x_{0}\sim p(x)$. The forward process is defined as:
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}),$ (13)
where $\beta_{t}\in(0,1)$ are predefined variance schedule parameters controlling the noise level at each time step $t\in[1,2,...,T]$, and $\mathbf{I}$ is the identity matrix.
For simplicity, we define $\alpha_{t}=1-\beta_{t}$, and denote the cumulative product $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. The reverse process, which is learned by the model $\theta$, can be defined as:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}\left(x_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right),\Sigma_{\theta}(x_{t},t)\right),$ (14)
where $\epsilon_{\theta}(x_{t},t)$ denotes the predicted noise from a neural network. The covariance $\Sigma_{\theta}(x_{t},t)$ is typically set to $\sigma^{2}_{t}\mathbf{I}$, where $\sigma^{2}_{t}$ can be either fixed ($\sigma^{2}_{t}=\beta_{t}$) or learned through interpolation $\sigma^{2}_{t}=(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})\beta$.
A simplified training objective minimizes the prediction error between true and estimated noise:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\|^{2}\right]$ (15)
In addition to the simple objective, improved variants include learning the reverse variance $\Sigma_{\theta}(x_{t},t)$ jointly with the mean, which leads to a variational bound loss of the form:
$\mathcal{L}_{\text{vlb}}=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right)$ (16)
Here, $v$ is an element-wise weight across model output dimensions. When $T$ is sufficiently large and the noise schedule is carefully chosen, the terminal distribution $p(x_{T})$ approximates an isotropic Gaussian. Sampling is then performed by iteratively applying the learned reverse process to recover the data sample from pure noise.
B.2 Diffusion Transformer Architecture
Our model implementation closely follows the design of DiT [11] and SiT [13], which extend the vision transformer (ViT) architecture [41] to generative modeling. An input image is first split into patches, reshaped into a 1D sequence of length $N$, and then processed through transformer layers. To reduce spatial resolution and computational cost, we follow prior work [11, 13] and encode the image into a latent tensor $z=E(x)$ using a pretrained encoder $E$ from the stable diffusion VAE [42].
In contrast to the standard ViT, our transformer blocks include time-aware adaptive normalization layers known as AdaIN-zero. These layers scale and shift the hidden state in each attention block according to the diffusion timestep and conditioning signals. During training, we also add an auxiliary multilayer perceptron (MLP) head that maps the hidden state to a semantic target representation space, such as DINOv2 [36] or CLIP features [43]. This head is used only for training-time supervision in our alignment loss and does not affect sampling or inference.
Appendix C Hyperparameters and Implementation Details
Select Phase Settings. In the Select phase, we adopt a pre-trained DiT-XL/2 model [11] as the scoring network and use the diffusion loss (mean squared error) as the scoring metric. To construct subsets of different sizes, we apply interval sampling with $k=96$ for the 10K subset, $k=16$ for the 50K subset, and $k=10$ for the 100K subset. Each subset is constructed in a class-wise manner, selecting 10, 50, and 100 samples per class respectively.
Attach Phase Settings. In the Attach phase, we implement dual conditional embeddings. For textual conditioning, we use a T5 encoder [35] with captions truncated to 16 tokens, producing embeddings of dimension 2048. For visual conditioning, we adopt DINOv2-B [36] as the visual encoder. The number of visual tokens $h$ is set to 256, and each token has a feature dimension of 768.
Training Settings. In the Training phase, we use the Adam optimizer with a fixed learning rate of 1e-4 and $(\beta_{1},\beta_{2})=(0.9,0.999)$, without applying weight decay. We employ mixed-precision (fp16) training with gradient clipping. Latent representations are pre-computed using the stable diffusion VAE [42], and decoded via its native decoder. All experiments are conducted on either 8 NVIDIA A800 80GB GPUs or 8 NVIDIA RTX 4090 24GB GPUs. We use a batch size of 256 with a $256\times 256$ resolution in Figure 1, and a $512\times 512$ resolution in Table 2. All other experiments use a batch size of 128 and a default image resolution of $256\times 256$.
Appendix D Framework Design and Implementation
The design of D2C aims to construct compact yet effective training subsets for diffusion models under strict data budgets. It is built on two complementary intuitions: (1) not all training samples contribute equally—some are easier to learn and more informative; and (2) generative training benefits from semantically enriched conditioning. These insights motivate the Select phase, which ranks samples by diffusion-estimated difficulty, using a pre-trained class-conditional diffusion model to compute a scoring function over training examples. The Attach phase then injects textual and visual priors. While empirical in nature, the framework draws from ideas in curriculum learning, semantic conditioning, and representation alignment. The full process is summarized in Algorithm 1.
Algorithm 1 D2C: Diffusion Dataset Condensation
0:  Full dataset $\mathcal{D}=\{(x_{i},c_{i})\}_{i=1}^{N}$, interval $k$, text encoder $f_{\text{text}}$, visual encoder $f_{\text{vis}}$ // Each $x_{i}$ is an image, and $c_{i}\in\{1,\dots,C\}$ is the class label.
1:  // Phase 1: Select
2:  Compute difficulty score $s_{\text{diff}}$ for all $(x_{i},c_{i})\in\mathcal{D}$
3:  For each class $c$, sort $\mathcal{D}_{c}=\{x_{i}\mid c_{i}=c\}$ by $s_{\text{diff}}$ descending
4:  Select every $k$-th sample (Interval Sampling) in sorted $\mathcal{D}_{c}$ to form $\mathcal{D}_{\text{select}}$
5:  // Phase 2: Attach
6:  for each $(x,c)\in\mathcal{D}_{\text{select}}$ do
7:     Generate class prompt $P(c)$ (e.g., “a photo of a label”)
8:     Extract text embedding: $(t_{c},t_{\textrm{mask}})\leftarrow f_{\text{text}}(P(c))$
9:     Extract visual feature: $y_{\text{vis}}\leftarrow f_{\text{vis}}(x)$
10:     Store triplet $(x,c,t_{c},t_{\textrm{mask}},y_{\text{vis}})$ into $\widetilde{\mathcal{D}}$
11:  end for
12:  Return enriched dataset $\widetilde{\mathcal{D}}$ for diffusion model training
Appendix E More Discussions about Select
Figure 8: Left: gFID-10K across training steps under different interval values $k$ for a 50K data budget. Moderate intervals (e.g., $k=16$) achieve superior performance by balancing learnability and diversity. Right: Distributional discrepancy (gFID-10K) between ranked training subsets and the validation set. Both extremely low and high diffusion difficulty scores lead to higher FID, while mid-range segments show better alignment.
While Section 4.2 has covered a detailed ablation study on the choice of interval $k$ in Select phase, we provide additional insights into how diffusion difficulty scores relate to distributional coverage.
The right panel in Fig. 8 presents the gFID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset. We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using gFiD. Interestingly, we observe a clear U-shaped curve: subsets consisting of extremely low or high difficulty samples exhibit significantly worse distributional alignment, while those centered around moderate difficulty levels show substantially lower FID scores. This result aligns well with our hypothesis that very easy samples (e.g., simple textures, clean backgrounds) and extremely hard samples (e.g., ambiguous, noisy structures) both fail to reflect the global data distribution.
These observations provide an empirical justification for our interval sampling strategy. Specifically, under a 50K dataset budget with $k=16$, each class contributes samples selected at regular intervals from its difficulty-sorted list. Given that each class typically contains around 1,200 images, this strategy naturally samples from approximately the first 800 positions in the ranked list. As a result, the selected data span both the easy and moderately difficult regions, while avoiding the extremes at both ends. This balanced coverage across the difficulty spectrum promotes better generalization and faster convergence, as evidenced by the results in Fig. 8 (Left) and discussed in Section 4.2. In this way, our strategy yields a compact yet effective dataset that enables the model to converge rapidly while maintaining strong generation quality.
Figure 9: t-SNE visualization of class embeddings. Each point represents a class in the dataset. Left: One-hot class embeddings show no semantic structure. Right: Text embeddings naturally cluster semantically related classes.
Appendix F More Discussion about Attach
F.1 Dual Conditional Embedding
Most diffusion models condition on class identifiers represented as integer IDs or one-hot vectors, which are mapped to class embeddings trained from scratch. This ignores semantic relationships between categories, resulting in unstructured embeddings as shown in Figure 9 (Left).In contrast, text embeddings derived from class-specific prompts (e.g., “a photo of a dog”) via a pre-trained language encoder naturally encode semantic priors and cluster related classes (Figure 9, Right). We propose a dual conditional embedding that fuses the text embedding with a learnable class embedding (i.e., a traditional class token trained from scratch), as defined in Eq. 89. This hybrid strategy combines semantic structure with symbolic distinctiveness, and leads to significantly improved generation quality. As shown in Figure 6 (Right), using both branches achieves lower FID than using either one alone.
F.2 Visual Information Injection
Recent studies[12] have shown that relying solely on diffusion models to learn meaningful representations from scratch often results in suboptimal semantic features. In contrast, injecting high-quality visual priors, especially those derived from strong self-supervised encoders like DINOv2[36], can significantly improve both training efficiency and generation quality. In our case, we incorporate a frozen visual encoder (DINOv2) to provide external patch-level visual features during training. These external features serve as semantically rich anchors, particularly beneficial at early layers, allowing the model to focus on generation-specific details in later stages. Empirically, we observe that such visual supervision leads to better feature alignment and faster convergence under limited data budgets, as evidenced in Table 12, and 4.
Appendix G Evaluation Details
We adopt several widely used metrics to evaluate generation quality and diversity:
  • gFID [38] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [44].
  • sFID [45] extends FID by leveraging intermediate spatial features from the Inception-v3 model to better capture spatial structure and style in generated images.
  • IS [39] evaluates both the quality and diversity of generated samples by computing the KL-divergence between the conditional label distribution and the marginal distribution over predicted classes, using softmax-normalized logits.
  • Precision and Recall [46] respectively measure sample realism and diversity, quantifying how well generated samples cover the data manifold and vice versa.
Appendix H Baseline Setting
We evaluate our method against two categories of baselines:
Diffusion models trained on selected or condensed subsets. These include SiT and DiT backbones trained from scratch on 10K, 50K, and 100K subsets obtained via the following strategies:
  • Random Sampling. A naive baseline that randomly selects a fixed number of real samples without any guidance.
  • Herding [31]. A geometry-based method that selects samples to approximate the global feature mean, ensuring representative coverage.
  • K-Center [30]. A diversity-focused algorithm that iteratively selects samples maximizing the minimum distance from the selected set, promoting broad coverage of the feature space.
  • SRe2L [24]. A dataset condensation method that synthesizes class-conditional data through a multi-stage pipeline. Originally proposed for classification tasks, we adapt it to the diffusion setting by applying class-wise condensation to real images and training a diffusion model on the resulting synthetic subset. Visualizations of the synthesized samples and corresponding training results are provided in Appendix I.
Diffusion models trained on the full dataset. These baselines are trained with access to the entire training set, without data reduction:
  • SiT [13]. A transformer-based diffusion model that reformulates denoising as continuous stochastic interpolation, enabling faster training and improved efficiency under full-data settings.
  • REPA [12]. A model-side regularization method that aligns intermediate features of diffusion transformers with patch-wise representations from strong pretrained visual encoders (e.g., DINOv2-B [36], MAE [47], MoCov3 [48]) using a contrastive loss. It retains the full dataset and improves convergence and generation quality via early-layer representation guidance.
Appendix I Visualization of SRe2L in Generative Tasks
Figure 10: Top: images synthesized directly by SRe2L, a popular dataset condensation method originally designed for discriminative tasks. Bottom: images generated by a diffusion model trained on the SRe2L dataset. As exemplified by SRe2L, such methods often struggle in generative settings—producing blurry, low-fidelity outputs that are poorly aligned with the true data distribution.
Appendix J ImageNet 512$\times$512 Experiment
As shown in Table 5, D2C consistently outperforms random sampling under a strict 10K (0.8%) data budget across both DiT-L/2 and SiT-L/2 backbones. At 300k iterations, our method achieves a gFID of 5.8 on DiT-L/2 and 4.22 on SiT-L/2, showing strong improvements. Visual samples in Figure 11 further confirm the high fidelity and diversity of generations at 512$\times$512 resolution, demonstrating that D2C generalizes effectively to high-resolution settings.
Table 5: Comparison with a data budget 0.8% (10K) on ImageNet 512$\times$512 (CFG=1.5).
[…]
DiT-L/2 Random 100k 24.8 11.9 74.3 0.65 0.42
DiT-L/2 D2C (Ours) 100k 14.8 6.9 109.2 0.63 0.52
DiT-L/2 Random 300k 17.1 12.8 130.6 0.64 0.41
DiT-L/2 D2C (Ours) 300k 5.8 15.1 318.9 0.77 0.29
SiT-L/2 Random 100k 13.3 22.8 197.1 0.69 0.68
SiT-L/2 D2C (Ours) 100k 9.1 14.3 261.7 0.72 0.34
SiT-L/2 Random 300k 5.0 13.6 316.9 0.76 0.27
SiT-L/2 D2C (Ours) 300k 4.22 11.6 289.7 0.79 0.24
Figure 11: Generated samples on ImageNet 512$\times$512 from SiT-L/2 trained with D2C using a 10K dataset (CFG=1.5).
Appendix K Visualization
Figure 12: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "macaw"(88)
Figure 13: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "arctic wolf"(270)
Figure 14: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "jaguar"(290)
Figure 15: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "otter"(360)
Figure 16: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "lesser panda"(387)
Figure 17: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "panda"(388)
Figure 18: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "fire truck"(555)
Figure 19: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "cheeseburger"(933)
Figure 20: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "lake shore"(975)
Figure 21: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "volcano"(980)

### v2

@@
 This unified training strategy enables D2C to effectively learn from limited yet enhanced data, offering a practical solution for efficient diffusion training under resource-constrained settings.
-Table 1: Comparison of gFID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet 256$\times$256.
+Table 1: Comparison of FID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet 256$\times$256.
 We use CFG=1.5 for evaluation.
@@
 However, due to the limited diversity of the selected dataset, performance tends to be overtaken by random sampling in later iterations.
-As the interval value increases, the final gFID-10K scores initially improve and then decline, suggesting an optimal interval value exists.
+As the interval value increases, the final FID-10K scores initially improve and then decline, suggesting an optimal interval value exists.
 Notably, this optimal value is 96 for a 10K data budget and 16 for a 50K data budget, and the ratio 96/16 closely approximates the ratio of the data budgets, 50K/10K.
@@
 Evaluation and baselines.
-We train models from scratch on the collected subset and evaluate them using gFID [38], sFID, Inception Score [39], Precision (Prec.), and Recall (Rec.), adhering to standard evaluation protocols [16, 11, 13].
+We train models from scratch on the collected subset and evaluate them using FID [38], sFID, Inception Score [39], Precision (Prec.), and Recall (Rec.), adhering to standard evaluation protocols [16, 11, 13].
 We compare our method against REPA [12], and various data condensation and selection baselines, including SRe2[24], Herding, K-Center, and random sampling, using SiT and DiT architectures [13, 11].
@@
 As illustrated in Fig.
-1 (b), our method achieves a gFID-50K of 4.23 at only 40K iterations with 10K training data and without CFG.
+1 (b), our method achieves a FID-50K of 4.23 at only 40K iterations with 10K training data and without CFG.
 In contrast, REPA requires 4 million steps and the vanilla SiT model needs over 7 million steps to reach comparable performance, representing an acceleration of over 100$\times$ and 233$\times$, respectively.
@@
 As illustrated in Table 1, D2C consistently achieves the lowest FID across all settings.
-For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a gFID-50K of 4.20 on DiT-L/2 and 3.98 on SiT.
+For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a FID-50K of 4.20 on DiT-L/2 and 3.98 on SiT.
 These results demonstrate the superiority of our approach over existing methods.
@@
 Comparison on ImageNet 512$\times$512.
-As shown in Table 2, D2C achieves a gFID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet 512$\times$512 settings.
+As shown in Table 2, D2C achieves a FID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet 512$\times$512 settings.
 On SiT-L/2, similar improvements are observed.
@@
 3).
-As $k$ increases, we observe that gFID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability.
+As $k$ increases, we observe that FID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability.
 Empirically, the best results are achieved with an interval of 96 for the 10K budget and 16 for 50K, approximately following the ratio of data budgets (50K/10K).
@@
 Second, as shown in Table 4, we analyze the impact of the semantic and visual information injection modules.
-The gFID-10K increases to 37.07 when both branches are removed and to 9.01 when only the visual injection module is excluded.
+The FID-10K increases to 37.07 when both branches are removed and to 9.01 when only the visual injection module is excluded.
 Notably, the combination of both modules yields the best FID of 7.62.
@@
 Appendix E More Discussions about Select
-Figure 8: Left: gFID-10K across training steps under different interval values $k$ for a 50K data budget.
+Figure 8: Left: FID-10K across training steps under different interval values $k$ for a 50K data budget.
 Moderate intervals (e.g., $k=16$) achieve superior performance by balancing learnability and diversity.
-Right: Distributional discrepancy (gFID-10K) between ranked training subsets and the validation set.
+Right: Distributional discrepancy (FID-10K) between ranked training subsets and the validation set.
 Both extremely low and high diffusion difficulty scores lead to higher FID, while mid-range segments show better alignment.
@@
 The right panel in Fig.
-8 presents the gFID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset.
-We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using gFiD.
+8 presents the FID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset.
+We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using FID.
 Interestingly, we observe a clear U-shaped curve: subsets consisting of extremely low or high difficulty samples exhibit significantly worse distributional alignment, while those centered around moderate difficulty levels show substantially lower FID scores.
@@
 We adopt several widely used metrics to evaluate generation quality and diversity:
-  • gFID [38] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [44].
+  • FID [38] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [44].
   • sFID [45] extends FID by leveraging intermediate spatial features from the Inception-v3 model to better capture spatial structure and style in generated images.
@@
 As shown in Table 5, D2C consistently outperforms random sampling under a strict 10K (0.8%) data budget across both DiT-L/2 and SiT-L/2 backbones.
-At 300k iterations, our method achieves a gFID of 5.8 on DiT-L/2 and 4.22 on SiT-L/2, showing strong improvements.
+At 300k iterations, our method achieves a FID of 5.8 on DiT-L/2 and 4.22 on SiT-L/2, showing strong improvements.
 Visual samples in Figure 11 further confirm the high fidelity and diversity of generations at 512$\times$512 resolution, demonstrating that D2C generalizes effectively to high-resolution settings.

### v3

@@
-Diffusion Dataset Condensation: Training Your Diffusion Model Faster with Less Data
+Accelerating Diffusion Model Training under Minimal Budgets: A Condensation-Based Perspective
 Abstract
-Diffusion models have achieved remarkable success in various generative tasks, but training them remains highly resource-intensive, often with millions of images and GPU days of computation required.
-From a data-centric perspective addressing the limitation, we study diffusion dataset condensation as a new challenging problem setting that aims at constructing a “synthetic” sub-dataset with significantly fewer samples than the original dataset for training high-quality diffusion models significantly faster.
-To the best of our knowledge, we are the first to formally study the dataset condensation task for diffusion models, while conventional dataset condensation focused on training discriminative models.
-For this new challenge, we further propose a novel Diffusion Dataset Condensation (D2C) framework, that consists of two phases: Select and Attach.
-The Select phase identifies a compact and diverse subset via a diffusion difficulty score and interval sampling, upon which the Attach phase enhances conditional signals and information of the selected subset by attaching rich semantic and visual representations.
-Extensive experiments across dataset sizes, model architectures, and resolutions demonstrate that our D2C can train diffusion models significantly faster with dramatically fewer data while retaining high visual quality.
-Notably, for the SiT-XL/2 architecture, our D2C achieves a 100$\times$ acceleration, reaching a FID of 4.3 in just 40k steps using only 0.8% of the training data.
+[…]
 footnotetext: Equal contribution.
    Correspondence to zekexie@hkust-gz.edu.cn.
-Figure 1: Our D2C framework significantly accelerates diffusion model training with limited data.
-(a) Overview of our D2C pipeline, which consists of a Select phase that filters a compact and diverse subset via diffusion difficulty score and interval sampling, and an Attach phase that enriches samples with semantic and visual information.
-(b) D2C achieves over 100$\times$ faster convergence compared to REPA and over 233$\times$ faster than vanilla SiT-XL/2, reaching a FID of 4.3 at just 40k steps.
-(c) Under a strict 4% data budget (0.05M), our method achieves a FID of 2.7 at 180k iterations, demonstrating its strong training efficiency and rapid convergence.
+[…]
 1 Introduction
-Generative models, such as score-based [1, 2, 3, 4] and flow-based [5, 6, 7] approaches, have achieved remarkable success in various generative tasks, producing high-quality and diverse data across domains [8, 9, 10].
+Generative models, such as score-based [43, 42, 17] and flow-based [27] approaches, have achieved remarkable success in various generative tasks [3], producing high-quality and diverse data across domains [19, 12, 4].
 However, these approaches are notoriously data and compute intensive to train, often requiring millions of samples and hundreds of thousands of iterations to capture complex high-dimensional distributions [33, 53, 28].
-The resulting cost presents a significant barrier to broader application and iteration within the AIGC community, making efficient training increasingly important across both academic and industrial settings [14, 15].
-Recent efforts have improved diffusion training efficiency through various strategies, such as architectural redesigns [13, 16, 11], attention optimization [17], reweighting strategies [18, 19, 20, 8], and representation learning [12].
-In parallel, data-centric approaches such as Patch Diffusion [21] and advanced augmentation techniques [8] aim to better exploit the potential of existing data.
-Despite these advancements, the exploration of building a relatively complete “synthetic” subset through dataset condensation [22] to accelerate diffusion model training with minimal information loss remains underexplored.
-Dataset condensation [22, 23, 24, 25] aims to construct a “synthetic” sub-dataset with significantly fewer samples than the original dataset, such that a model trained from scratch on this subset achieves comparable performance to one trained on the full dataset.
-Unlike data pruning or selection [26], which passively select existing samples, condensation actively optimizes synthetic data, offering greater potential for aggressive data reduction and training efficiency [27].
-However, all existing methods are designed for discriminative tasks.
-Compared to discriminative tasks, generative tasks are much more complex and demand higher dataset quality [28].
-Applying popular methods (e.g., SRe2[24]) that have substantiated effective for discriminative tasks to diffusion models presents significant challenges, such as the failure to produce diverse, high-quality outputs with structural and semantic fidelity, leading to degraded results and unstable convergence (see Sec.
+The resulting cost presents a significant barrier to broader application and iteration within the AIGC community, making efficient training increasingly important across both academic and industrial settings [41, 26, 11].
+Recent efforts have improved diffusion training efficiency through various strategies, such as architectural redesigns [28, 55, 33], attention optimization [1, 24], reweighting strategies [13], and representation learning [53, 48, 54].
+In parallel, data-centric approaches such as patch-based methods [7, 47], Infobatch [34] and Reweighting[25] aim to better exploit the potential of existing data.
+Despite these advances, directly constructing a much smaller yet informative training subset as a primary condensation-based route to accelerate diffusion training remains largely unexplored, even though it provides a particularly direct and effective way to reduce data budgets.
+Dataset condensation [49, 46, 52, 38] aims to construct a much smaller condensed sub-dataset with significantly fewer samples than the original dataset, such that a model trained from scratch on this subset achieves performance comparable to one trained on the full dataset while converging much faster.
+In practice, existing DC methods typically instantiate this objective in two ways [49]: (i) pixel-level dataset distillation, which directly optimizes synthetic images [52, 2], and (ii) image-level condensation, which operates on real images through selection and transformation [44, 10, 20].
+Unlike classical data pruning or selection [51], which passively select a fixed subset of existing samples, both branches of dataset condensation actively construct condensed training sets, either by optimizing synthetic images or by selecting and enriching real ones, thereby enabling more aggressive data reduction and higher training efficiency [44].
+However, these methods have been developed almost exclusively for discriminative tasks.
+Compared to discriminative learning, generative diffusion training is substantially more complex and demands higher dataset quality [29]; directly applying popular DC algorithms (e.g., SRe2[52], RDED [44]) to diffusion models often leads to synthetic images that lack structural and semantic fidelity, resulting in degraded sample quality and unstable convergence (see Sec.
 4).
 We raise a key question: “Can we train diffusion models dramatically faster with significantly less data, while retaining high generation quality?” The answer is affirmative.
-Answering this question holds significant relevance for the further development of visually generative intelligence and is therefore extremely worthwhile to explore.
-In this paper, we mainly made the following contributions.
-First, to the best of our knowledge, we are the first to formally study the dataset condensation task for diffusion models, a new challenging problem setting that aims at constructing a “synthetic” sub-dataset with significantly fewer samples than the original dataset for training high-quality diffusion models significantly faster.
+In this paper, we make three main contributions.
+First, to the best of our knowledge, we are the first to formally study the dataset condensation task for diffusion models, a new challenging problem setting that aims at constructing a “condensed” sub-dataset with significantly fewer samples than the original dataset for training high-quality diffusion models significantly faster.
+We address a fundamental academic gap concerning the application of dataset condensation in diffusion models.
 More specifically, our explorations with the diffusion model provide the first insights into the challenges and potential solutions for applying dataset condensation to vision generation tasks.
-We note that while conventional dataset condensation made great progress and sometimes use diffusion models to construct a subset, this line of research only focused on training discriminative models instead of generative models.
-Second, we propose a a novel two-stage Diffusion Dataset Condensation (D2C) framework, consisting of Select and Attach, illustrated in Fig.
-1 (a).
+We note that while conventional dataset condensation has made great progress and sometimes uses diffusion models to construct a subset, this line of research only focused on training discriminative models instead of generative models.
+Second, we propose D2C, a novel two-stage dataset condensation framework tailored for training diffusion models.
 Our framework addresses the challenges of dataset condensation for diffusion models by decomposing the problem into two key aspects: the Select stage identifies an informative, compact, and learnable subset by ranking samples using the diffusion difficulty score derived from a pre-trained diffusion model; the Attach stage enriches each selected sample by adding semantic and visual representations, further enhancing the training efficiency while preserving performance.
-Third, extensive experiments demonstrate that the proposed D2C can train diffusion models significantly faster with dramatically fewer data while retaining high visual quality, substantiating the effectiveness and scalability.
-Specifically, D2C significantly outperforms random sampling and several popular dataset selection and distillation algorithms (e.g., SRe2[24] and K-Center) across data compression ratios of 0.8%, 4%, and 8%, at resolutions of 256$\times$256 and 512$\times$512, and with both SiT [13] and DiT [11] architectures.
-In particular, D2C achieves a FID of 4.3 in merely 40k training steps (w/o classifier-free guidance (CFG) [4]) using SiT-XL/2 [13], demonstrating a 100$\times$ acceleration over REPA [12] and a 233$\times$ speed-up compared to vanilla SiT.
-Furthermore, it further improves to a FID of 2.7 using only 50k synthesized images with CFG (refer to Fig.
-1 (c)).
+[…]
 2 Preliminaries and Related Work
-In this section, we briefly review diffusion models as well as dataset condensation.
 Diffusion Models.
-We briefly introduce the standard latent-space noise injection formulation that underlies many diffusion models [6, 11, 9].
-These models define a forward process that gradually perturbs input data with Gaussian noise, and train a neural network to approximate the reverse denoising process for sampling (see Appendix B for more details).
-Let the original variable be denoted as $\mathbf{x}_{0}\sim q_{0}(\mathbf{x})$. The forward diffusion process is typically implemented as a Markov chain with Gaussian noise injections:
+We briefly introduce the standard latent-space noise injection formulation [33], which defines a forward process that gradually perturbs input data $\mathbf{x}_{0}\sim q_{0}(\mathbf{x})$ with Gaussian noise:
 $q_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}),$ (1)
@@
 The choice for $\alpha_{t}$ and $\sigma_{t}$ is referred to as the noise schedule of a diffusion model.
-A commonly used reparameterization in score-based diffusion models [3] expresses the noisy data $\mathbf{x}_{t}$ at time step $t$ as $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\epsilon$, where $\ \epsilon\sim\mathcal{N}(0,\mathbf{I})$.
-In this case, to recover “clean” data from noise, a neural network $\epsilon_{\theta}(\cdot,\cdot,\cdot)$ is trained to predict the added noise $\epsilon$, thereby approximating the reverse process.
-This framework adopt the conventional noise prediction strategy for training, where the training objective is to minimize the mean squared error between the predicted and the ground true noise:
+After that, we need to train a neural network $\epsilon_{\theta}(\cdot,\cdot,\cdot)$ to approximate the reverse denoising process (i.e., predict the added noise $\epsilon$) for sampling (see Appendix B for more details).
+The training objective is to minimize the mean squared error between the predicted and the ground true noise:
 $\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\|^{2}_{2}\right],$ (2)
-Here, $\mathbf{c}$ is a conditional input, such as class labels or text embeddings.
-In some cases, the prediction target is replaced with a linear combination of noise and data (e.g.
-$v$-prediction), which corresponds to flow matching [5, 6, 7]:
-$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-(\epsilon-\mathbf{x}_{0})\|^{2}_{2}\right].$ (3)
-Here, $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ denotes the learned time-dependent velocity field.
+Here, $\mathbf{c}$ is a conditional input, such as class labels or text embeddings. In some cases, the prediction target is replaced with the $v$-prediction, which corresponds to flow matching.
+Figure 2: Overview of Diffusion Dataset Condensation (D2C).
+D2C employs a two-stage process: Select and Attach.
+The Select stage identifies a compact and diverse subset by interval sampling using the diffusion difficulty score derived from a pre-trained diffusion model.
+The Attach stage further enriches each selected sample by adding semantic information and visual information.
 Data-centric Efficient Training.
-Various model-side strategies have been proposed to accelerate diffusion model, including architectural enhancements [17, 11, 13], sampling refinements [18, 19, 20, 8], and representation-level techniques that leverage pretrained vision features [12, 29].
-However, relatively little has been explored from a data-centirc perspective.
-In data-centric model training, given an original dataset $\mathcal{D}=(\hat{\mathbf{X}},\hat{\mathbf{Y}})=\{(\hat{\mathbf{x}}_{i},\hat{y}_{i})\}_{i=1}^{|\mathcal{D}|}$, where each $\hat{y}_{i}$ is the label corresponding to sample $\hat{\mathbf{x}}_{i}$, dataset compression aims to reduce the size of training data while preserving model performance.
+Beyond model-side improvements, a complementary line of work takes a data-centric view on diffusion efficiency.
+Patch-based schemes [7, 47] and Infobatch [34] focus on reallocating training effort over existing samples by reweighting or resampling informative regions and instances.
+However, comparatively few methods directly tackle diffusion training efficiency by explicitly reducing and restructuring the overall training set.
+In this setting, given an original dataset $\mathcal{D}=\{(\hat{\mathbf{x}}_{i},\hat{y}_{i})\}_{i=1}^{|\mathcal{D}|}$, where each $\hat{y}_{i}$ is the label corresponding to sample $\hat{\mathbf{x}}_{i}$, dataset compression aims to reduce the size of training data while preserving model performance.
 Two primary strategies have been extensively studied in this context: dataset pruning and dataset condensation.
 1) Dataset Pruning.
-Dataset pruning selects an information-enrichment subset from the original dataset, i.e., $\mathcal{D}^{\text{core}}\subset\mathcal{D}$ with $|\mathcal{D}^{\text{core}}|\ll|\mathcal{D}|$, and directly minimizes the training loss over the subset:
+Dataset pruning selects an information-enriched subset from the original dataset, i.e., $\mathcal{D}^{\text{core}}\subset\mathcal{D}$ with $|\mathcal{D}^{\text{core}}|\ll|\mathcal{D}|$, and directly minimizes the training loss over the subset:
 $\min_{\theta}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}^{\text{core}}}\left[\ell(\phi_{\theta_{\mathcal{D}^{\text{core}}}}(\mathbf{x}),y)\right],$ (3)
@@
 Classical data pruning methods like random sampling, K-Center [18], and Herding [5] can be used with diffusion models, but they offer minimal performance improvements.
+Very recently, Li et al.
+[25] investigate data-efficient diffusion training from the perspective of dataset pruning by selecting a coreset with surrogate features and then performing class-wise reweighting.
+While this approach substantially reduces training cost and improves over naive pruning, it does not attach any additional information to the selected samples and is mainly validated on relatively small-scale or latent diffusion settings, which limits its ability to fully exploit the potential of condensed training data for large-scale, high-resolution diffusion models.
 2) Dataset Condensation.
-In contrast, dataset condensation aims to synthesize a small, compact, and diverse synthetic dataset $\mathcal{D}^{\mathcal{S}}=(\mathbf{X},\mathbf{Y})=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{|\mathcal{D}^{\mathcal{S}}|}$ to replace the original dataset $\mathcal{D}$.
+Following recent work [49], dataset condensation aims to synthesize a small, compact, and diverse synthetic dataset $\mathcal{D}^{\mathcal{S}}=(\mathbf{X},\mathbf{Y})=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{|\mathcal{D}^{\mathcal{S}}|}$ to replace the original dataset $\mathcal{D}$.
 The synthetic dataset $\mathcal{D}^{\mathcal{S}}$ is generated by a condensation algorithm $\mathcal{C}$ such that $\mathcal{D}^{\mathcal{S}}\in\mathcal{C}(\mathcal{D})$, with $|\mathcal{D}^{\mathcal{S}}|\ll|\mathcal{D}|$.
@@
 $\sup\left\{\left|\ell(\phi_{\theta_{\mathcal{D}}}(\hat{\mathbf{x}}),\hat{y})-\ell(\phi_{\theta_{\mathcal{D}}^{\mathcal{S}}}(\hat{\mathbf{x}}),\hat{y})\right|\right\}_{(\hat{\mathbf{x}},\hat{y})\sim\mathcal{D}}\leq\eta,$ (4)
-where $\theta_{\mathcal{D}}$ is the parameter set of the neural network $\phi$ optimized on $\mathcal{D}$: $\theta_{\mathcal{D}}=\arg\min_{\theta}\mathbb{E}_{(\hat{\mathbf{x}},\hat{y})\sim\mathcal{D}}\left[\ell(\phi_{\theta}(\hat{\mathbf{x}}),\hat{y})\right].$ A similar definition applies to $\theta_{\mathcal{D}}^{\mathcal{S}}$, which is optimized on the synthetic dataset $\mathcal{D}^{\mathcal{S}}$.
-Existing methods, such as RDED, MTT, SRe2L, and G-VBSM [27, 24, 25, 32, 33], are primarily designed for discriminative tasks.
-When applied to diffusion models, these methods generate synthetic images that deviate from the original data distribution, leading to detrimental effects on model training.
-Visualization of these synthesized images can be found at Appendix I.
-Figure 2: Overview of Diffusion Dataset Condensation (D2C).
-D2C employs a two-stage process: Select and Attach.
-The Select stage identifies a compact and diverse subset by intervaling sampling using the diffusion difficulty score derived from a pre-trained diffusion model.
-The Attach stage further enriches each selected sample by adding semantic information and visual information.
+[…]
 3 Diffusion Dataset Condensation
-To enable data-centric efficient training of diffusion models under limited resources, we propose Diffusion Dataset Condensation (D2C), the first unified framework that systematically condenses training data for diffusion models.
 As illustrated in Fig.
-2, this process produces a condensed dataset suitable for efficient diffusion model training.
-D2C consists of two stages: Select (Sec.
+2, D2C consists of two stages: Select (Sec.
 3.1), which identifies a compact set of diverse and learnable real images using diffusion difficulty score and interval sampling techniques; and Attach (Sec.
 3.2), which augments each selected image with semantic and visual information to improve generation performance.
-Finally, we present the novel training paradigm utilizing the condensed dataset generated by D2C (Sec.
+Finally, we describe how to train diffusion models on the condensed dataset produced by D2C in Sec.
+3.3.
+(Sec.
 3.3).
 3.1 Select: Difficulty-Aware Selection
-In this work, we focus on class-to-image (C2I) synthesis.
-Given a class-conditioned dataset $\mathcal{D}=\bigcup_{y=1}^{C}\mathcal{D}_{y}$, where $\mathcal{D}_{y}=\{x_{i}\}_{i=1}^{|\mathcal{D}_{y}|}$ denotes all samples of class $y$, we aim to select a compact subset for efficient diffusion training.
-To achieve this, we propose diffusion difficulty score to quantify the denoising difficulty of each sample, followed by our designed interval sampling to ensure diversity within the selected subset.
+In this work, we focus on class-to-image (C2I) synthesis, aligned with the setting in Yu et al.
+[53], and show that our framework also applies to the text-to-image (T2I) setting with only minor changes; see Appendix G for details.
+Given a class-conditioned dataset $\mathcal{D}=\bigcup_{y=1}^{C}\mathcal{D}_{y}$, where $C$ denotes the class number and $\mathcal{D}_{y}=\{x_{i}\}_{i=1}^{|\mathcal{D}_{y}|}$ denotes all samples of class $y$, our aim is to select a compact subset for efficient diffusion training.
+To achieve this, we propose the diffusion difficulty score to quantify the denoising difficulty of each sample, followed by our designed interval sampling to ensure diversity within the selected subset.
 Diffusion Difficulty Score.
 The arrangement of samples from easy to hard is crucial for revealing underlying data patterns and facilitating difficulty-aware selection.
-Recent work [34] demonstrates that diffusion models inherently encode semantic-related class-conditional probability $p_{\theta}(\mathbf{c}|\mathbf{x})$ through the variational lower bound (i.e., diffusion loss Eq.
-2) of $\log p_{\theta}(\mathbf{x}|\mathbf{c})$ [3, 1].
-This conditional probability can be formulated as $p_{\theta}(\mathbf{c}|\mathbf{x})=\frac{p_{\theta}(\mathbf{x}|\mathbf{c})p(\mathbf{c})}{\sum_{\mathbf{\hat{c}}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})p(\mathbf{\hat{c}})}$.
-Intuitively, a larger $p_{\theta}(\mathbf{c}|\mathbf{x})$ indicates that sample $\mathbf{x}$ can be more confidently identified as belonging to class $\mathbf{c}$, thus suggesting lower learning difficulty.
-Given the significant computational overhead of the full Bayesian formulation and our focus on estimating sample difficulty, we ignore the denominator part $\sum_{\mathbf{\hat{c}}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})p(\mathbf{\hat{c}})$ of the calculation.
-Since the category $y\sim U\{0,\cdots,C\}$ ($C$ denotes the class number) is obtained by uniform sampling, and assuming $\sup\{|\mathbb{E}_{\mathbf{\hat{c}}}[p_{\theta}(\mathbf{x_{1}}|\mathbf{\hat{c}})]-|\mathbb{E}_{\mathbf{\hat{c}}}[p_{\theta}(\mathbf{x_{2}}|\mathbf{\hat{c}})]|\}_{\mathbf{x_{1}},\mathbf{x}_{2}\sim\mathcal{D^{X}}}\leq\eta$ ($\mathcal{D^{X}}$ denotes the all images in the dataset), we define the diffusion difficulty score based on the class-conditional probability $p_{\theta}(\mathbf{x}|\mathbf{c})\propto\frac{p_{\theta}(\mathbf{x}|\mathbf{c})}{\sum_{\mathbf{c}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})}=p_{\theta}(\mathbf{c}|\mathbf{x})$ ($\sum_{\mathbf{c}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})$ can be viewed as a constant when sorting by sample difficulty):
-$s_{\text{diff}}(\mathbf{x})=-p_{\theta}(\mathbf{x}|\mathbf{c})=-\log\left(\exp\left(-\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\left\|\epsilon-\epsilon_{\theta}\left(\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon,\,t,\,\mathbf{c}\right)\right\|^{2}_{2}\right]\right)\right),$ (6)
-The higher the score $s_{\text{diff}}(\mathbf{x})$, the easier it is, and the lower the score $s_{\text{diff}}(\mathbf{x})$, the more difficult it is.
-Note that in this paper’s figures and tables, we directly use the diffusion loss (i.e., $\propto-p_{\theta}(\mathbf{c}|\mathbf{x})$) for ease of understanding.
-By computing $s_{\text{diff}}(x)$ for all training samples, we construct a ranked dataset.
-As shown in Fig.
-3 (Left), these scores exhibit a skewed unimodal distribution.
-Selecting the easiest samples (Min) yields a subset dominated by clean, background-simple images with high learnability but limited diversity.
-In contrast, selecting only the highest-score samples (Max) results in cluttered, noisy, and ambiguous images that are difficult to optimize.
-Meanwhile, many samples lie in the middle range, offering moderate learnability but richer contextual information.
-Interval Sampling.
-To balance diversity and learnability, we propose an interval sampling strategy.
-Specifically, we sort its images $\mathcal{D}_{y}$ within each class $y$ in ascending order of $s_{\text{diff}}(x)$ and select samples at a fixed interval $k$:
-$\mathcal{D}_{\text{IS}}=\bigcup_{y=1}^{C}\left\{x^{(i)}\in\mathcal{D}_{y}\;\middle|\;i\in\{0,k,2k,\dots\}\right\},$ (7)
-where $\mathcal{D}_{\text{IS}}$ denotes the selected subset constructed by interval sampling, $k$ is the fixed sampling interval, and $x^{(i)}$ is the $i$-th sample in the sorted list (e.g., $x^{(0)}$ corresponds to the sample with the lowest diffusion difficulty score).
-Interval sampling with a larger interval $k$ promotes diversity in the sampled data while potentially hindering learnability.
-As shown in Fig.
-3 (Left), this trade-off arises from a shift in the sample distribution: a larger $k$ leads to a reduction in the number of easy samples and a corresponding increase in the representation of standard and difficult samples.
+Recent work [23, 56] demonstrates that diffusion models inherently encode semantic-related class-conditional probability $p_{\theta}(\mathbf{c}\mid\mathbf{x})$ through the variational lower bound (i.e., diffusion loss Eq.
+2) of $\log p_{\theta}(\mathbf{x}\mid\mathbf{c})$ [17, 43].
+This conditional probability admits the standard Bayesian form
+$p_{\theta}(\mathbf{c}\mid\mathbf{x})=\frac{p_{\theta}(\mathbf{x}\mid\mathbf{c})\,p(\mathbf{c})}{\sum_{\hat{c}}p_{\theta}(\mathbf{x}\mid\hat{c})\,p(\hat{c})}$. (5)
+Intuitively, a larger $p_{\theta}(\mathbf{c}\mid\mathbf{x})$ indicates that sample $\mathbf{x}$ can be more confidently identified as belonging to class $\mathbf{c}$, thus suggesting lower learning difficulty.
+Computing the full denominator in Eq.
+(5) for every sample is expensive, while we only need a score that orders samples by difficulty.
+Since the class label $y\sim U\{1,\dots,C\}$ is obtained by uniform sampling and the average likelihood over classes does not vary too much across samples, i.e., we assume $\sup_{\mathbf{x}_{1},\mathbf{x}_{2}\sim\mathcal{D}}\bigl|\mathbb{E}_{\hat{c}}[p_{\theta}(\mathbf{x}_{1}\mid\hat{c})]-\mathbb{E}_{\hat{c}}[p_{\theta}(\mathbf{x}_{2}\mid\hat{c})]\bigr|\leq\eta$,where $\eta>0$ is a small tolerance, the denominator in Eq.
+(5) can be treated as approximately constant with respect to $\mathbf{x}$.
+Consequently, the posterior is proportional to the class-conditional likelihood,
+$p_{\theta}(\mathbf{c}\mid\mathbf{x})\propto p_{\theta}(\mathbf{x}\mid\mathbf{c}).$ (6)
+We define the diffusion difficulty score based on this posterior:
+$s_{\text{diff}}(\mathbf{x})$ $=-p_{\theta}(\mathbf{c}\mid\mathbf{x})\propto-p_{\theta}(\mathbf{x}\mid\mathbf{c})$ (7)
+$=-\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I}),\,t\sim\mathcal{U}[0,1]}\left[\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{x_{t}},t,\mathbf{c}\right)\right\|^{2}_{2}\right]$
+The higher the score $s_{\text{diff}}(\mathbf{x})$, the more difficult it is, and the lower the score $s_{\text{diff}}(\mathbf{x})$, the easier it is.
+To simplify our presentation, we define the diffusion loss $p_{\theta}(\mathbf{x}|\mathbf{c})$ as the diffusion difficulty score.
 Figure 3: Left: Distribution of diffusion difficulty scores under different interval values $k$.
@@
 Interval sampling achieves a balance between structural clarity and contextual richness.
-Following interval sampling yields notable findings.
-First, interval sampling efficiently mitigates the inclusion of both overly simplistic and excessively challenging samples, yielding subsets that exhibit both learnability and diversity.
-Fig.
-3 (right) displays representative examples selected by our strategy, showcasing structural clarity coupled with contextual richness.
-Second, the FID of diffusion models trained on the subset $\mathcal{D}_{\text{IS}}$ generated with increasing $k$ initially decreases and subsequently increases.
-For example, as shown in Fig.
-8, our results further corroborate that $k=16$ strikes a favorable balance between sample diversity and generation quality when using a 50K data budget.
+By computing $s_{\text{diff}}(x)$ for all training samples, we construct a ranked dataset.
+As shown in Fig.
+3, these scores exhibit a skewed unimodal distribution.
+Selecting the easiest samples (Min) yields a subset dominated by clean, background-simple images with high learnability but limited diversity.
+In contrast, selecting only the highest-score samples (Max) results in cluttered, noisy, and ambiguous images that are difficult to optimize.
+Meanwhile, many samples lie in the middle range, offering moderate learnability but richer contextual information.
+Selecting an appropriate value within this range is therefore critical; we provide a more detailed discussion in Appendix H.1.
+Interval Sampling.
+To balance diversity and learnability, we propose an interval sampling strategy.
+Specifically, we sort its images $\mathcal{D}_{y}$ within each class $y$ in ascending order of $s_{\text{diff}}(x)$ and select samples at a fixed interval $k$: $\mathcal{D}_{\text{IS}}=\bigcup_{y=1}^{C}\left\{x^{(i)}\in\mathcal{D}_{y}\;\middle|\;i\in\{0,k,2k,\dots\}\right\}$, where $\mathcal{D}_{\text{IS}}$ denotes the selected subset constructed by interval sampling, $k$ is the fixed sampling interval, and $x^{(i)}$ is the $i$-th sample in the sorted list (e.g., $x^{(0)}$ corresponds to the sample with the lowest diffusion difficulty score).
+Interval sampling with a larger interval $k$ promotes diversity in the sampled data while potentially hindering learnability.
+As shown in Fig.
+3 (Left), this trade-off arises from a shift in the sample distribution: a larger $k$ leads to a reduction in the number of easy samples and a corresponding increase in the representation of standard and difficult samples.
+Figure 4: Overview of DC-Embedding.
+Extended Discussion.
+Training exclusively on the easiest (Min) or the hardest (Max) samples is suboptimal.
+Instead, a balanced curriculum comprising easy, medium, and difficult examples yields a training subset that is both learnable and diverse, ultimately leading to stronger generative performance.
+We further offer more discussions and insights on interval sampling in Appendix H.2.
 3.2 Attach: Semantic and Visual Information Enhancement
-To further enhance the information richness of the selected data, we introduce the Attach phase that augments each instance with semantic and visual information.
-While the Select phase generates a compact and informative subset, the achievable performance ceiling based solely on this phase remains limited.
-Consequently, we inject more comprehensive semantic and visual representations into the selected subset to further bolster our method’s generalization capability.
+To complement the Select phase, which yields a compact subset of informative real images, the Attach phase enriches each selected instance with additional semantic and visual information.
+In particular, we attach semantic information via a Dual Conditional Embedding (DC-Embedding) module and inject visual information through visual representation, resulting in a more expressive condensed dataset and improved generalization of the trained diffusion models.
 Dual Conditional Embedding (DC-Embedding).
-Existing C2I synthesis methods [11, 13] commonly rely on class embeddings trained from scratch, which often fail to effectively capture inherent semantic information (see Appendix F.1).
+Existing C2I synthesis methods [33, 28] commonly rely on class embeddings trained from scratch, which often fail to effectively capture inherent semantic information (see Appendix I.1).
 We enrich the class embedding by incorporating text representations derived from a pre-trained text encoder (e.g., T5-encoder [31]).
 For each class $c\in\{1,\dots,C\}$, a descriptive prompt $P(c)$ (e.g., “a photo of a cat”) is encoded by a pre-trained text encoder $f_{\text{text}}$, yielding its corresponding text embedding $t_{c}$ and text mask $t_{\textrm{mask}}$:
-$\begin{split}t_{c},t_{\textrm{mask}}&=f_{\text{text}}(P(c)),\\ \end{split}$ (8)
+[…]
 The resulting text embedding and text mask are stored on disk as attached text information alongside the subset $\mathcal{D}_{\textrm{IS}}$ generated in the preceding phase, ready for import during formal training.
 During the formal training, as illustrated in Fig.
-5, the text embedding $t_{c}$ and the text mask $t_{\textrm{mask}}$ undergo a 1D convolution and are fused with a learnable class embedding $e_{c}$ using a residual MLP:
-[…]
-Figure 4: Visualization of samples from Random (i.e., random sampling) vs.
-D2C.
-D2C demonstrates a significant performance advantage over Random across all training iterations, under both a strict 0.8% (10K) and a 4% (50K) data budget, leading to a substantial acceleration in training.
-Figure 5: Overview of DC-Embedding.
+4, the text embedding $t_{c}$ and the text mask $t_{\textrm{mask}}$ undergo a 1D convolution and are fused with a learnable class embedding $e_{c}$ using a residual MLP:
+$\tilde{t}_{c}$ $=\text{Conv1d}(t_{c}\times t_{\textrm{mask}}),\quad y_{\text{text}}=\text{MLP}(\tilde{t}_{c})+\tilde{t}_{c}+e_{c}.$ (9)
+Figure 5: D2C improves visual quality under tight data budgets. We compare Random sampling and D2C on DiT-L/2 at 10k and 50k data budgets, and neither setting uses classifier-free guidance.
 This resulting vector $y_{\text{text}}$ then serves as a semantic conditioning token for the conditional diffusion model.
@@
 $y_{\text{vis}}=f_{\text{vis}}(x)\in\mathbb{R}^{N\times d_{\text{text}}}$ (10)
-[…]
+where $N$ is the number of image patches and $d_{\text{text}}$ is the feature dimension.
+We retain the first $h$ (i.e., number of tokens in the diffusion transformer) tokens of $y_{\text{vis}}$ to form a compact representation of the dominant structure: $y_{\text{vis}}=y_{\text{vis}}[:\!h,:]\in\mathbb{R}^{h\times d_{\text{text}}}$.
+As outlined in REPA [53], this visual information provides a semantic prior for the diffusion model and thus significantly benefits data-centric efficient training.
+Similar to the text information $y_{\text{text}}$, the visual information $y_{\text{vis}}$ is also stored on disk as attached metadata alongside the selected subset $\mathcal{D}_{\text{IS}}$.
 3.3 D2C Training Process
 Here, we detail the training process of the diffusion model using our condensed dataset, which comprises a compact subset selected during the Select phase and subsequently enriched with semantic and visual information during the Attach phase.
-Our goal is to fully leverage the information existed within our condensed dataset to accelerate training without compromising performance.
+Our goal is to fully leverage the information contained in our condensed dataset to accelerate training without compromising performance.
 We employ a conditional diffusion model $\mathcal{D}_{\theta}$ and, as an example, utilize the optimization objective of score-based diffusion models: predicting the added noise $\epsilon$ from the perturbed latent input $\mathbf{x}_{t}$ at time step $t$, conditioned on the text information $y_{\textrm{text}}$ and the class label $y$.
-The new denoising loss is defined as $\mathcal{L}^{\text{new}}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,y,y_{\textrm{text}})\|^{2}_{2}\right]$, where the specific injected forms of $y$ and $y_{\text{text}}$ can be found in Sec.
+The new denoising loss is defined as $\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,y,y_{\textrm{text}})\|^{2}_{2}\right]$, where the specific injected forms of $y$ and $y_{\text{text}}$ can be found in Sec.
 3.2.
@@
 $\mathcal{L}_{\text{proj}}=-\frac{1}{h}\sum_{i=1}^{h}\left\langle\frac{\phi(h_{i})}{\|\phi(h_{i})\|},\frac{v_{i}}{\|v_{i}\|}\right\rangle.$ (11)
-This loss encourages the model to align its encoder’s output with visual representations extracted from the visual encoder, promoting localized realism and spatial consistency [36] in generation.
+This loss encourages the model to align its encoder’s output with visual representations, promoting localized realism and spatial consistency [32] in generation.
 Overall Training Objective. The final training loss combines the denoising objective and the semantic alignment term (with the balance weight $\lambda$ is set to 0.5 by default):
 $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{diff}}+\lambda\mathbb{E}_{\mathbf{x},\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1],y,y_{\text{text}},y_{\text{vis}}}\left[\mathcal{L}_{\text{proj}}\right].$ (12)
-This unified training strategy enables D2C to effectively learn from limited yet enhanced data, offering a practical solution for efficient diffusion training under resource-constrained settings.
-Table 1: Comparison of FID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet 256$\times$256.
+This training strategy enables D2C to effectively learn from limited yet enhanced data, offering a practical solution for efficient diffusion training under minimal budgets.
+4 Experiments
+In this section, we validate the performance of D2C and analyze the contributions of its components through extensive experiments.
+In particular, we aim to answer the following questions: 1) Can D2C improve training speed and reduce data usage of diffusion models?
+2) Does D2C generalize well across backbones, data scales, and resolutions?
+3) How do D2C’s components and hyperparameter choices affect its overall effectiveness?
+Table 1: Comparison of gFID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet $\times256$.
 We use CFG=1.5 for evaluation.
@@
 8.0% (100K) 300k 11.49 37.35 15.23 6.49 12.56 39.08 16.17 5.65
-Table 2: Comparison with a strict data budget 0.8% (10K) on ImageNet 512$\times$512.
-We use CFG=1.5 for evaluation.
-D2C surpasses random sampling at all settings.
-D2C’s performance on SiT-L/2 with 300k iterations can be found in Appendix J.
-[…]
-DiT-L/2 Random 100k 24.8 11.9 74.3 0.65 0.42
-DiT-L/2 D2C (Ours) 100k 14.8 6.9 109.2 0.63 0.52
-DiT-L/2 Random 300k 17.1 12.8 130.6 0.64 0.41
-DiT-L/2 D2C (Ours) 300k 5.8 15.1 318.9 0.77 0.29
-SiT-L/2 Random 100k 13.3 22.8 197.1 0.69 0.68
-SiT-L/2 D2C (Ours) 100k 9.1 14.3 261.7 0.72 0.34
-4 Experiments
-We validate the performance of D2C and analyze the contributions of its components through extensive experiments. In particular, we aim to answer the following questions:
-1) Can D2C improve training speed and reduce data usage of diffusion models? (Table 1, Figs. 14)
-2) Does D2C generalize well across backbones, data scales, and resolutions? (Tables 12, Fig. 4)
-3) How do D2C’s components and hyperparameter choices affect its overall effectiveness? (Figs. 368)
+4.1 Setup
 Experiment settings.
 We conduct experiments on the ImageNet-1K dataset [36], using subsets of 10K, 50K, and 100K images, corresponding to 0.8%, 4%, and 8% of the full dataset, respectively.
-All images are center-cropped and resized to 256$\times$256 and 512$\times$512 resolutions using the ADM [16] preprocessing pipeline.
-Furthermore, we use [$\cdot$]-L/2 and [$\cdot$]-XL/2 architectures in both DiT [11] and SiT [13] backbones, following the standard settings outlined in Ma et al.
+To further demonstrate the generalization and effectiveness of our method, Appendix J reports additional results of D2C on CIFAR datasets.
+All images are center-cropped and resized to $\times256$ and $\times512$ resolutions using the ADM [6] preprocessing pipeline.
+Furthermore, we use $\cdot$]-L/2 and $\cdot$]-XL/2 architectures in both DiT [33] and SiT [28] backbones, following the standard settings outlined in Ma et al.
 [28].
-More details on implementation and training can be found in Appendix C.
-Figure 6: Left: Ablation studies of the interval value $k$ in interval sampling at different data budgets.
-With a small interval value, training can experience significant acceleration during the early iterations.
-However, due to the limited diversity of the selected dataset, performance tends to be overtaken by random sampling in later iterations.
-As the interval value increases, the final FID-10K scores initially improve and then decline, suggesting an optimal interval value exists.
-Notably, this optimal value is 96 for a 10K data budget and 16 for a 50K data budget, and the ratio 96/16 closely approximates the ratio of the data budgets, 50K/10K.
-Right:.
-Ablation study of DC-Embedding under a 10K data budget.
-“Only Class” stands for the baseline approach of injecting class embeddings.
-This figure demonstrates that both the text embedding branch and the original class embedding branch (see Fig.
-5) are critical for the effective training of diffusion models.
-Table 3: D2C vs. SRe2[24] on ImageNet 256$\times$256 with a strict data budget 0.8% (10K).
+Table 2: Comparison with a strict data budget 0.8% (10K) on ImageNet $\times512$. We use CFG=1.5 for evaluation. D2C surpasses random sampling at all settings.
 […]
-DiT-L/2 SRe2L 104.2 20.2 14.1 0.20 0.30
-DiT-L/2 D2C (Ours) 4.2 11.0 283.6 0.72 0.24
-SiT-L/2 SRe2L 82.3 19.8 18.1 0.27 0.19
-SiT-L/2 D2C (Ours) 3.9 10.7 289.7 0.73 0.25
+DiT-L/2 Random 100k 24.8 11.9 74.3 0.65
+DiT-L/2 D2C (Ours) 100k 14.8 6.9 109.2 0.63
+DiT-L/2 Random 300k 17.1 12.8 130.6 0.64
+DiT-L/2 D2C (Ours) 300k 5.8 15.1 318.9 0.77
+SiT-L/2 Random 100k 13.3 22.8 197.1 0.69
+SiT-L/2 D2C (Ours) 100k 9.1 14.3 261.7 0.72
+SiT-L/2 Random 300k 5.0 13.6 316.9 0.76
+SiT-L/2 D2C (Ours) 300k 4.22 11.6 289.7 0.79
 Evaluation and baselines.
-We train models from scratch on the collected subset and evaluate them using FID [38], sFID, Inception Score [39], Precision (Prec.), and Recall (Rec.), adhering to standard evaluation protocols [16, 11, 13].
-We compare our method against REPA [12], and various data condensation and selection baselines, including SRe2[24], Herding, K-Center, and random sampling, using SiT and DiT architectures [13, 11].
-Further details regarding evaluation metrics and baseline methods can be found in Appendix G and H.
+We train models from scratch on the collected subset and evaluate them using gFID [16], sFID, Inception Score [37] and Precision, adhering to standard evaluation protocols [6, 33, 28].
+We compare our method against REPA [53], REPA-E [22], REG [48] and various data condensation and selection baselines, including SRe2[52], RDED [44], Herding, K-Center, and random sampling, using SiT and DiT architectures [28, 33].
+Further details regarding evaluation metrics and baseline methods can be found in Appendix D and E.
+Figure 6: Left: Interval-sampling ablation.
+Small $k$ speeds early training.
+The best final gFID-10K appears at k=96k{=}96 for the 10K budget and k=16k{=}16 for the 50K budget, roughly scaling with data size.
+Right: DC-Embedding ablation at 10K.
+Combining text and class embeddings outperforms either alone; “Only Class” denotes the baseline that injects class embeddings only.
 4.2 Main Result
-Training Performance and Speed.
-We evaluate D2C on SiT-XL/2 using 10K and 50K data budgets, comparing its performance against REPA and a vanilla SiT model trained on the full ImageNet dataset (a 1.28M data budget), as well as random selection with 10K and 50K data budgets.
-As illustrated in Fig.
-1 (b), our method achieves a FID-50K of 4.23 at only 40K iterations with 10K training data and without CFG.
-In contrast, REPA requires 4 million steps and the vanilla SiT model needs over 7 million steps to reach comparable performance, representing an acceleration of over 100$\times$ and 233$\times$, respectively.
-Under a 4% data budget (50K) with CFG set to 1.5, our method achieves an FID of 2.78 at 180K steps, further demonstrating significant data and compute efficiency (Fig.
-1 (c)).
-Moreover, Fig.
-4 presents a visual comparison between random selection and our D2C at 10K and 50K data sizes, where images are generated from the same noise and class label.
-Our method demonstrates superior visual quality compared to the baseline and generates higher-quality images, even during the early iterations of training.
-Comparison on ImageNet 256$\times$256.
-We compare D2C with random sampling, Herding [40], K-Center [30], and SRe2[24] under various data budgets and backbones.
-As illustrated in Table 1, D2C consistently achieves the lowest FID across all settings.
-For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a FID-50K of 4.20 on DiT-L/2 and 3.98 on SiT.
+[…]
+Table 3: Comparison of acceleration algorithms on ImageNet-1K.
+[…]
+DiT L-2 1.28M 400k 23.3
+++ REPA 1.28M 400k 15.6
+++ D2C 0.05M 10k 14.81
+++ D2C 0.01M 10k 4.2
+SiT L-2 1.28M 400k 18.8
+++ REPA 1.28M 700k 8.4
+++ D2C 0.01M 80k 7.07
+SiT XL-2 1.28M 7M 8.3
+++ REPA 1.28M 4M 5.9
+++ REPA-E 1.28M 235k 5.9
+++ REG 1.28M 200k 5.0
+++ D2C 0.01M 40k 4.3
+++ D2C 0.05M 180k 2.78
+Comparison on ImageNet $\times256$.
+We compare D2C with random sampling, Herding [5], K-Center [18], SRe2[52], RDED [44] under various data budgets and backbones.
+As shown in Table 1, D2C consistently achieves the lowest FID across all settings.
+For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a gFID-50K of 4.20 on DiT-L/2 and 3.98 on SiT.
 These results demonstrate the superiority of our approach over existing methods.
-Notably, SRe2L, which performs well in classification task, fails on this generative task (see Table 3) due to its focus on category-discriminative features.
-Similarly, geometry-based methods like Herding and K-Center, along with random sampling, prove inadequate for achieving efficient and high-performing training of diffusion models.
-Comparison on ImageNet 512$\times$512.
-As shown in Table 2, D2C achieves a FID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet 512$\times$512 settings.
+Notably, SRe2L and RDED, which perform well in classification task, fail on this generative task (see Table 4) due to their focus on category-discriminative features.
+Similarly, geometry-based methods like Herding and K-Center, along with random sampling, prove inadequate for achieving efficient and high-performing training.
+Comparison on ImageNet $\times512$.
+As shown in Table 2, D2C achieves a gFID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet $\times512$ settings.
 On SiT-L/2, similar improvements are observed.
-These results demonstrate that D2C generalizes well to higher resolutions.
+These demonstrate that D2C generalizes well to higher resolutions.
 4.3 Ablation Study
-Figure 7: Ablation sutdies of interval sampling and DC-Embedding. “Mean=N” denotes the selection of a subset of samples where the average diffusion loss is closest to the value N.
+Table 4: D2C vs. SRe2[52] and RDED [44] on ImageNet $\times256$ with a data budget 0.8% (10K).
+[…]
+DiT-L/2 RDED 166.2 60.1 10.8 0.09
+DiT-L/2 SRe2L 104.2 20.2 14.1 0.20
+DiT-L/2 D2C (Ours) 4.2 11.0 283.6 0.72
+SiT-L/2 RDED 97.5 66.8 65.63 0.22
+SiT-L/2 SRe2L 82.3 19.8 18.1 0.27
+SiT-L/2 D2C (Ours) 3.9 10.7 289.7 0.73
 Ablation on Select Phase.
 We investigate the impact of the interval value $k$ in the Select phase, as shown in Fig.
-6 (Left) and Fig.
-7.
+6 (Left).
 Using a small value accelerates early training by prioritizing min-loss samples, which are simpler and easier to learn.
@@
 3).
-As $k$ increases, we observe that FID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability.
+As $k$ increases, we observe that gFID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability.
 Empirically, the best results are achieved with an interval of 96 for the 10K budget and 16 for 50K, approximately following the ratio of data budgets (50K/10K).
+Table 5 further shows that using the Select stage alone reduces gFID from 37.07 to 14.96, underscoring its effectiveness and usefulness.
+Ablation on Attach Phase.
+We evaluate Attach from two angles.
+First, as shown in Fig.
+6 (Right), DC embedding consistently outperforms using either alone under a 10K budget, with text-only better than class-only, indicating richer semantics from textual descriptions.
+Second, Table 5 shows steady gains from the injection modules: baseline gFID-10K is 14.96, adding only visual information reaches 10.37, adding only DC embedding reaches 9.01, and combining both achieves the best 7.62.
+Appendix I.2 further ablates the visual encoder and demonstrates the robustness of our approach.
+Table 5: Ablation studies on the Select and Attach phases. Sel.: Select. Vis.: Vision.
 […]
 DiT-L/2 37.07
+DiT-L/2 8.79
+DiT-L/2 14.96
+DiT-L/2 10.37
 DiT-L/2 9.01
 DiT-L/2 7.62
-Table 4: Ablation on semantic and visual information in the Attach phase.
-Ablation on Attach Phase.
-We conduct ablation studies on the Attach phase from two perspectives.
-First, as shown in Fig.
-6 (Right), we examine the effectiveness of dual conditional embedding under a 10K data budget.
-Using only class or text embedding results in degraded performance compared to their combination, indicating the necessity of both components to capture category semantics and text embeddings.
-Notably, text embedding alone performs better than class embedding, suggesting that textual descriptions provide richer semantic information for generative tasks.
-Second, as shown in Table 4, we analyze the impact of the semantic and visual information injection modules.
-The FID-10K increases to 37.07 when both branches are removed and to 9.01 when only the visual injection module is excluded.
-Notably, the combination of both modules yields the best FID of 7.62.
-These results underscore the crucial role of injecting both semantic and visual information in accelerating diffusion model training.
+Table 6: A breakdown of the computational overhead for sub-processes in D2C. Compared to the REPA baseline, the additional scoring time is negligible, demonstrating D2C’s efficiency.
+[…]
+REPA N/A N/A 4M 750h 5.9
+D2C (w/o select) N/A N/A 0.04M 7.4h 5.6
+D2C (w/ select) From Scratch 1.9h 0.04M (7.4+26.2)h 4.9
+D2C (w/ select) Pretrained 2.1h 0.04M 7.4h 4.3
+[…]
 5 Conclusion
-In this paper, we introduce D2C, the first dataset condensation franework that can significantly accelerate diffusion model training for the generative task.
-Our pipeline comprises two key phases: Select and Attach.
-In the Select phase, we leverage diffusion difficulty score and interval sampling to obtain a subset that is both compact and diverse.
-Subsequently, in the Attach phase, we augment this subset with critical semantic and visual information, leading to a dramatic improvement in the training acceleration ratio and enhanced performance robustness.
-While there are still some limitations (refer to Appendix A), our D2C as the pioneering work in this novel approach still achieved impressively 100233×100$\sim$ 233$\times$100 ∼ 233 × training acceleration over the baselines.
-We believe that our work will offer novel inspirations and motivate more research on this promising field in the future.
-Broader Impact
-We introduce D2C, a method engineered to decrease training overhead while minimizing performance degradation.
-The dataset employed for training D2C was ImageNet-1K, which adheres to ethical guidelines.
-Moreover, we prioritize the responsible and ethical deployment of this technology, aiming to maximize its societal benefits while proactively addressing and mitigating any potential risks.
-Appendix
-Appendix A Limitation
-This work primarily focuses on accelerating and improving the training process of diffusion models.
-While our approach demonstrates promising results in the early stage, the final performance upon convergence does not yet fully match baseline benchmarks.
-It is important to unleash the potential of our algorithm, such as fine-tuning the vision and text encoders to better align noise latent features with the integrated visual information, thereby reducing existing discrepancies.
-Additionally, although the current study centers on image generation tasks, extending our method to other domains, such as 3D generation and video synthesis, represents a promising direction for future exploration.
+In this paper, we introduce D2C, the first dataset condensation framework that significantly accelerates diffusion model training for generative tasks.
+D2C follows a two-phase pipeline, Select and Attach, which selects a compact yet diverse subset via a diffusion difficulty score with interval sampling and enriches it with semantic and visual signals.
+On ImageNet-1K, D2C achieves $100\text{--}233\times$ faster training than strong baselines while maintaining competitive generative quality, and we hope it will motivate further research on data-centric efficiency for diffusion models.
+Acknowledgement. This work was supported by the National Natural Science Foundation of China under Grant No. 62506317.
+Supplementary Material
+Appendix A Positioning D2C within Dataset Condensation Paradigms
+While some works equate dataset condensation with gradient-based pixel-level optimization of synthetic images, a broader line of literature defines it as constructing compact training sets that retain the learning efficacy of the original data [49], which also includes image-level schemes such as OD3[20] and RDED[44].
+In this broader paradigm, the key objective is not how the condensed data are obtained, but whether the resulting small dataset can support training models that closely match the performance of those trained on the full dataset.
+D2C follows this latter view.
+It condenses the dataset by selecting a highly informative subset guided by diffusion difficulty and then attaching additional semantic and visual information that enriches each sample without altering its raw pixels.
+This design is analogous in spirit to OD3 and RDED, which also operate at the level of image selection rather than direct pixel optimization.
+Consequently, D2C naturally fits within the dataset condensation family, while being specifically tailored to generative diffusion models and addressing a gap that is not covered by existing pixel-level condensation methods.
 Appendix B Additional Descriptions of Diffusion Models
-This section presents the preliminaries of the diffusion model.
-For simplicity, we introduce the standard denoising diffusion probabilistic model (DDPM), where the forward process gradually corrupts the input with noise and the reverse process restores the target data distribution.
-We also summarize the architectural details in B.2.
-B.1 Denoising Diffusion Probabilistic Model (DDPM)
+This section reviews the fundamentals of the Denoising Diffusion Probabilistic Model (DDPM) [17].
+The DDPM framework consists of a fixed forward process that incrementally perturbs the input data with noise, and a learned reverse process trained to iteratively denoise the data, thereby learning the target distribution.
+Specific architectural details of our implementation are summarized in Appendix B.2.
+B.1 Denoising Diffusion Probabilistic Model
 The DDPM framework models data generation via a discrete-time Markov chain that progressively adds Gaussian noise to a data sample $x_{0}\sim p(x)$. The forward process is defined as:
@@
 The reverse process, which is learned by the model $\theta$, can be defined as:
-$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}\left(x_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right),\Sigma_{\theta}(x_{t},t)\right),$ (14)
+$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}\left(x_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right),\Sigma_{\theta}(x_{t},t)\right),$
 where $\epsilon_{\theta}(x_{t},t)$ denotes the predicted noise from a neural network.
@@
 A simplified training objective minimizes the prediction error between true and estimated noise:
-$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\|^{2}\right]$ (15)
+$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\|^{2}\right].$ (15)
 In addition to the simple objective, improved variants include learning the reverse variance $\Sigma_{\theta}(x_{t},t)$ jointly with the mean, which leads to a variational bound loss of the form:
-$\mathcal{L}_{\text{vlb}}=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right)$ (16)
+$\mathcal{L}_{\text{vlb}}=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right).$ (16)
 Here, $v$ is an element-wise weight across model output dimensions.
@@
 An input image is first split into patches, reshaped into a 1D sequence of length $N$, and then processed through transformer layers.
-To reduce spatial resolution and computational cost, we follow prior work [11, 13] and encode the image into a latent tensor $z=E(x)$ using a pretrained encoder $E$ from the stable diffusion VAE [42].
-In contrast to the standard ViT, our transformer blocks include time-aware adaptive normalization layers known as AdaIN-zero.
+To reduce spatial resolution and computational cost, we follow prior work [33, 28] and encode the image into a latent tensor $z=E(x)$ using a pretrained encoder $E$ from the stable diffusion VAE.
+In contrast to the standard ViT, our transformer blocks include time-aware adaptive normalization layers known as adaLN-zero.
 These layers scale and shift the hidden state in each attention block according to the diffusion timestep and conditioning signals.
@@
 Select Phase Settings.
-In the Select phase, we adopt a pre-trained DiT-XL/2 model [11] as the scoring network and use the diffusion loss (mean squared error) as the scoring metric.
+In the Select phase, we adopt a pre-trained DiT-XL/2 model [33] as the scoring network and use the diffusion loss (w.r.t., mean squared error) as the scoring metric.
 To construct subsets of different sizes, we apply interval sampling with $k=96$ for the 10K subset, $k=16$ for the 50K subset, and $k=10$ for the 100K subset.
@@
 We employ mixed-precision (fp16) training with gradient clipping.
-Latent representations are pre-computed using the stable diffusion VAE [42], and decoded via its native decoder.
+Latent representations are pre-computed using the stable diffusion VAE, and decoded via its native decoder.
 All experiments are conducted on either 8 NVIDIA A800 80GB GPUs or 8 NVIDIA RTX 4090 24GB GPUs.
-We use a batch size of 256 with a $256\times 256$ resolution in Figure 1, and a $512\times 512$ resolution in Table 2.
+We use a batch size of 256 with a $256\times 256$ resolution in Fig.
+1, and a $512\times 512$ resolution in Table 2.
 All other experiments use a batch size of 128 and a default image resolution of $256\times 256$.
-Appendix D Framework Design and Implementation
-The design of D2C aims to construct compact yet effective training subsets for diffusion models under strict data budgets.
-It is built on two complementary intuitions: (1) not all training samples contribute equally—some are easier to learn and more informative; and (2) generative training benefits from semantically enriched conditioning.
-These insights motivate the Select phase, which ranks samples by diffusion-estimated difficulty, using a pre-trained class-conditional diffusion model to compute a scoring function over training examples.
-The Attach phase then injects textual and visual priors.
-While empirical in nature, the framework draws from ideas in curriculum learning, semantic conditioning, and representation alignment.
-The full process is summarized in Algorithm 1.
+Appendix D Evaluation Details
+We adopt several widely used metrics to evaluate generation quality and diversity:
+  • gFID [16] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [45].
+  • sFID [30] extends FID by leveraging intermediate spatial features from the Inception-v3 model to better capture spatial structure and style in generated images.
+  • IS [37] evaluates both the quality and diversity of generated samples by computing the KL-divergence between the conditional label distribution and the marginal distribution over predicted classes, using softmax-normalized logits.
+  • Precision and Recall [21] respectively measure sample realism and diversity, quantifying how well generated samples cover the data manifold and vice versa.
+Appendix E Baseline Setting
+We evaluate our method against two categories of baselines:
+Diffusion models trained on selected or condensed subsets. These include SiT and DiT backbones trained from scratch on 10K, 50K, and 100K subsets obtained via the following strategies:
+  • Random Sampling. A naive baseline that randomly selects a fixed number of real samples without any guidance.
+  • Herding [5]. A geometry-based method that selects samples to approximate the global feature mean, ensuring representative coverage.
+  • K-Center [18]. A diversity-focused algorithm that iteratively selects samples maximizing the minimum distance from the selected set, promoting broad coverage of the feature space.
+  • SRe2L [52].
+A dataset condensation method that synthesizes class-conditional data through a multi-stage pipeline.
+Originally proposed for classification tasks, we adapt it to the diffusion setting by applying class-wise condensation to real images and training a diffusion model on the resulting synthetic subset.
+Visualizations of the synthesized samples and corresponding training results are provided in Appendix K.
+Diffusion models trained on the full dataset. These baselines are trained with access to the entire training set, without data reduction:
+  • SiT [28]. A transformer-based diffusion model that reformulates denoising as continuous stochastic interpolation, enabling faster training and improved efficiency under full-data settings.
+  • REPA [53].
+A model-side regularization method that aligns intermediate features of diffusion transformers with patch-wise representations from strong pretrained visual encoders (e.g., DINOv2-B [32], MAE [14], MoCov3 [15]) using a contrastive loss.
+It retains the full dataset and improves convergence and generation quality via early-layer representation guidance.
+Appendix F Framework Design and Implementation
+We introduce D2C, a framework for constructing compact yet effective training subsets for diffusion models under stringent data budgets.
+Our approach is motivated by two complementary intuitions: (1) that the contribution of training samples is non-uniform, as some are more informative than others; and (2) that generative training benefits from semantically enriched conditioning.
+These insights directly inform the two core stages of our framework.
+First, a Select stage ranks training examples by a difficulty score computed via a pretrained class-conditional diffusion model.
+Second, an Attach stage enriches the selected data by injecting textual and visual priors.
+The complete pipeline is summarized in Algorithm 1.
 Algorithm 1 D2C: Diffusion Dataset Condensation
@@
 2:  Compute difficulty score $s_{\text{diff}}$ for all $(x_{i},c_{i})\in\mathcal{D}$
-3:  For each class $c$, sort $\mathcal{D}_{c}=\{x_{i}\mid c_{i}=c\}$ by $s_{\text{diff}}$ descending
+3:  For each class $c$, sort $\mathcal{D}_{c}=\{x_{i}\mid c_{i}=c\}$ by $s_{\text{diff}}$ ascending
 4:  Select every $k$-th sample (Interval Sampling) in sorted $\mathcal{D}_{c}$ to form $\mathcal{D}_{\text{select}}$
@@
 12:  Return enriched dataset $\widetilde{\mathcal{D}}$ for diffusion model training
-Appendix E More Discussions about Select
-Figure 8: Left: FID-10K across training steps under different interval values $k$ for a 50K data budget.
+Appendix G Exploration on Text-to-Image Generation
+Figure 7: Distribution of diffusion difficulty score computed on LAION text–image pairs with a pre-trained SDXL model.
+This distribution resembles that of C2I, which supports interval sampling for selecting informative training pairs under T2I.
+We further examine the applicability of the D2C framework to text-to-image generation [50, 40, 9].
+The Select phase requires only a minimal change: replace the class condition in Eq.
+7 with a text condition, i.e., $s_{\mathrm{diff}}^{text}(x)=-p_{\theta}\!\left(x\mid\text{text}\right)$.
+Using SDXL to score LAION text–image pairs, we observe a difficulty distribution similar to the class-conditional case (Fig.
+7; see also Fig.
+3 and Fig.
+8 (right)).
+Low-score samples tend to exhibit simple structures, high-score samples often contain complex or cluttered contexts, and the majority of samples fall in the middle range.
+Interval sampling remains effective for identifying informative pairs.
+The Attach phase is also easy to transfer: semantic and visual representations serve as soft supervisory signals for the selected subset.
+As such, while our main experiments focus on class-to-image tasks for controlled benchmarking like SiT [28], the framework is generalizable and well suited to text-to-image generation.
+We expect it to deliver practical gains in data efficiency and training speed in this setting, offering a promising direction for future work.
+Appendix H More Discussions about Select
+H.1 Detailed Algorithm for Computing Diffusion Difficulty Score
+The diffusion difficulty score, used to rank samples in the Select phase, is defined as the mean denoising loss over uniformly sampled timesteps, computed using a frozen pretrained diffusion model (see Algorithm 2).
+Algorithm 2 Compute Diffusion Difficulty Score
+0:  Image dataset $\mathcal{D}=\{(x_{i},c_{i})\}_{i=1}^{N}$; pretrained VAE encoder $E_{\phi}$; pretrained diffusion model $\epsilon_{\theta}$; timestep set $\mathcal{T}$; batch size $n$ // Each $x_{i}$ is an image; $c_{i}\in\{1,\dots,C\}$ is the class label.
+Timesteps in $\mathcal{T}$ are sampled uniformly.
+Models are frozen during scoring.
+1:  Initialize empty map $\mathcal{S}\leftarrow\{\}$
+2:  for mini-batch $\{(x_{i},c_{i})\}_{i=1}^{n}\subset\mathcal{D}$ do
+3:   Encode to latent (if applicable): $z_{i}\leftarrow E_{\phi}(x_{i})$
+4:   Initialize per-sample accumulator $\ell_{i}\leftarrow 0$
+5:   for $t\in\mathcal{T}$ do
+6:    Sample $\epsilon\sim\mathcal{N}(0,I)$
+7:    Perturb latent: $z_{t}\leftarrow\alpha_{t}\,z_{i}+\sigma_{t}\,\epsilon$
+8:    Compute loss: $\ell_{i}\leftarrow\ell_{i}+\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c_{i})\rVert_{2}^{2}$
+9:   end for
+10:   $s_{i}\leftarrow\ell_{i}/|\mathcal{T}|$ // Mean denoising loss across timesteps
+11:   $\mathcal{S}[x_{i}]\leftarrow s_{i}$
+12:  end for
+13:  Return $\mathcal{S}$ // Image-to-score mapping for difficulty-aware selection
+H.2 Practical Insights on Interval Sampling
+Figure 8: Left: gFID-10K across training steps under different interval values $k$ for a 50K data budget.
 Moderate intervals (e.g., $k=16$) achieve superior performance by balancing learnability and diversity.
-Right: Distributional discrepancy (FID-10K) between ranked training subsets and the validation set.
-Both extremely low and high diffusion difficulty scores lead to higher FID, while mid-range segments show better alignment.
-While Section 4.2 has covered a detailed ablation study on the choice of interval $k$ in Select phase, we provide additional insights into how diffusion difficulty scores relate to distributional coverage.
+Right: Distributional discrepancy (gFID-10K) between ranked training subsets and the validation set.
+Both extremely low and high diffusion difficulty score lead to higher FID, while mid-range segments show better alignment.
+While Section 4.3 has covered a detailed ablation study on the choice of interval $k$ in Select phase, we provide additional insights into how diffusion difficulty score relates to distributional coverage.
 The right panel in Fig.
-8 presents the FID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset.
-We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using FID.
+8 presents the gFID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset.
+We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using gFID.
 Interestingly, we observe a clear U-shaped curve: subsets consisting of extremely low or high difficulty samples exhibit significantly worse distributional alignment, while those centered around moderate difficulty levels show substantially lower FID scores.
@@
 In this way, our strategy yields a compact yet effective dataset that enables the model to converge rapidly while maintaining strong generation quality.
+Ablation on interval sampling.
+As shown in Fig.
+9, the “Medium” variant corresponds to selecting samples from the center of the difficulty-ranked list rather than applying interval sampling from low to high diffusion difficulty scores.
+Concretely, after sorting each class by diffusion difficulty, we start from the median position and expand symmetrically toward both sides until the data budget is reached.
+This strategy focuses on medium-difficulty examples and largely omits easier instances, while still including a portion of harder ones near the tails.
+As a result, the selected subset provides less comprehensive coverage of the underlying data distribution, leading to slower convergence and degraded final performance compared to our proposed interval sampling scheme.
+Appendix I More Discussions about Attach
+I.1 Dual Conditional Embedding
 Figure 9: T-SNE visualization of class embeddings.
@@
 Right: Text embeddings naturally cluster semantically related classes.
-Appendix F More Discussion about Attach
-F.1 Dual Conditional Embedding
+Samples from semantically related classes, such as different dog breeds, tend to form distinct clusters in feature space.
+Leveraging this semantic prior is highly effective for accelerating diffusion model training.
 Most diffusion models condition on class identifiers represented as integer IDs or one-hot vectors, which are mapped to class embeddings trained from scratch.
-This ignores semantic relationships between categories, resulting in unstructured embeddings as shown in Figure 9 (Left).In contrast, text embeddings derived from class-specific prompts (e.g., “a photo of a dog”) via a pre-trained language encoder naturally encode semantic priors and cluster related classes (Figure 9, Right).
+This ignores semantic relationships between categories, resulting in unstructured embeddings as shown in Fig.
+9 (Left).In contrast, text embeddings derived from class-specific prompts (e.g., “a photo of a dog”) via a pre-trained language encoder naturally encode semantic priors and cluster related classes (Fig.
+9, Right).
 We propose a dual conditional embedding that fuses the text embedding with a learnable class embedding (i.e., a traditional class token trained from scratch), as defined in Eq.
@@
 This hybrid strategy combines semantic structure with symbolic distinctiveness, and leads to significantly improved generation quality.
-As shown in Figure 6 (Right), using both branches achieves lower FID than using either one alone.
-F.2 Visual Information Injection
-Recent studies[12] have shown that relying solely on diffusion models to learn meaningful representations from scratch often results in suboptimal semantic features.
+As shown in Fig.
+6 (Right), using both branches achieves lower FID than using either one alone.
+Figure 10: Top: Images synthesized directly by SRe2L and RDED, two popular dataset condensation methods originally designed for discriminative tasks.
+Bottom: Images generated by diffusion model trained on the two synthesized datasets.
+I.2 Visual Information Injection
+Table 7: Ablation of the visual encoder.
+[…]
+N/A (baseline) 37.07
+MAE-L 9.23
+MoCov3-L 8.78
+CLIP-L 8.59
+DINOv2-L 7.62
+Recent studies [48, 53] have shown that relying solely on diffusion models to learn meaningful representations from scratch often results in suboptimal semantic features.
 In contrast, injecting high-quality visual priors, especially those derived from strong self-supervised encoders like DINOv2 [32], can significantly improve both training efficiency and generation quality.
-In our case, we incorporate a frozen visual encoder (DINOv2) to provide external patch-level visual features during training.
+In our case, we incorporate a frozen visual encoder to provide external patch-level visual features during training.
 These external features serve as semantically rich anchors, particularly beneficial at early layers, allowing the model to focus on generation-specific details in later stages.
-Empirically, we observe that such visual supervision leads to better feature alignment and faster convergence under limited data budgets, as evidenced in Table 12, and 4.
-Appendix G Evaluation Details
-We adopt several widely used metrics to evaluate generation quality and diversity:
-  • FID [38] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [44].
-  • sFID [45] extends FID by leveraging intermediate spatial features from the Inception-v3 model to better capture spatial structure and style in generated images.
-  • IS [39] evaluates both the quality and diversity of generated samples by computing the KL-divergence between the conditional label distribution and the marginal distribution over predicted classes, using softmax-normalized logits.
-  • Precision and Recall [46] respectively measure sample realism and diversity, quantifying how well generated samples cover the data manifold and vice versa.
-Appendix H Baseline Setting
-We evaluate our method against two categories of baselines:
-Diffusion models trained on selected or condensed subsets. These include SiT and DiT backbones trained from scratch on 10K, 50K, and 100K subsets obtained via the following strategies:
-  • Random Sampling. A naive baseline that randomly selects a fixed number of real samples without any guidance.
-  • Herding [31]. A geometry-based method that selects samples to approximate the global feature mean, ensuring representative coverage.
-  • K-Center [30]. A diversity-focused algorithm that iteratively selects samples maximizing the minimum distance from the selected set, promoting broad coverage of the feature space.
-  • SRe2L [24].
-A dataset condensation method that synthesizes class-conditional data through a multi-stage pipeline.
-Originally proposed for classification tasks, we adapt it to the diffusion setting by applying class-wise condensation to real images and training a diffusion model on the resulting synthetic subset.
-Visualizations of the synthesized samples and corresponding training results are provided in Appendix I.
-Diffusion models trained on the full dataset. These baselines are trained with access to the entire training set, without data reduction:
-  • SiT [13]. A transformer-based diffusion model that reformulates denoising as continuous stochastic interpolation, enabling faster training and improved efficiency under full-data settings.
-  • REPA [12].
-A model-side regularization method that aligns intermediate features of diffusion transformers with patch-wise representations from strong pretrained visual encoders (e.g., DINOv2-B [36], MAE [47], MoCov3 [48]) using a contrastive loss.
-It retains the full dataset and improves convergence and generation quality via early-layer representation guidance.
-Appendix I Visualization of SRe2L in Generative Tasks
-Figure 10: Top: images synthesized directly by SRe2L, a popular dataset condensation method originally designed for discriminative tasks.
-Bottom: images generated by a diffusion model trained on the SRe2L dataset.
-As exemplified by SRe2L, such methods often struggle in generative settings—producing blurry, low-fidelity outputs that are poorly aligned with the true data distribution.
-Appendix J ImageNet 512$\times$512 Experiment
+Empirically, visual supervision improves feature alignment and accelerates convergence under limited data, as shown in Tables 1, 2, 5, and 7.
+All tested encoders outperform the no-encoder baseline, indicating that our method is robust to the choice of visual encoder.
+Appendix J Experiments on CIFAR
+As shown in Table 8, we further evaluate D2C on CIFAR-10 by selecting 100 images per class to form a 1K data budget (2% compression rate) and training the diffusion model for 100k steps.
+Under this highly constrained setting, D2C significantly improves gFID from 9.72 with random sampling to 3.95, demonstrating that our selection and attachment strategy remains effective beyond ImageNet and transfers well across datasets.
+Table 8: Comparison of random subset selection and D2C on CIFAR-10 (reported in gFID-50K).
+[…]
+Random 9.72
+D2C (Ours) 3.95
+Figure 11: Generated samples on ImageNet $\times512$ from SiT-L/2 trained with D2C using a 10K dataset (CFG=1.5).
+Appendix K Visualization of SRe2L and RDED in Generative Tasks
+As shown in Fig.
+10, dataset condensation methods that excel in classification, such as RDED and SRe2L, transfer poorly to diffusion-based generation.
+Their objectives focus on preserving class-discriminative cues, for example segmentation-guided selection in RDED and gradient-based image optimization in SRe2L, rather than modeling realistic global structure and natural image statistics.
+As a result, diffusion models trained on these synthesized datasets fail to capture the underlying pixel-level data distribution and produce severely degraded samples.
+In contrast, D2C provides the first dataset condensation framework tailored to diffusion generative modeling and effectively closes this gap.
+Appendix L ImageNet $\times512$ Experiment
 As shown in Table 2, D2C consistently outperforms random sampling under a strict 10K (0.8%) data budget across both DiT-L/2 and SiT-L/2 backbones.
-At 300k iterations, our method achieves a FID of 5.8 on DiT-L/2 and 4.22 on SiT-L/2, showing strong improvements.
-Visual samples in Figure 11 further confirm the high fidelity and diversity of generations at 512$\times$512 resolution, demonstrating that D2C generalizes effectively to high-resolution settings.
-Table 5: Comparison with a data budget 0.8% (10K) on ImageNet 512$\times$512 (CFG=1.5).
-[…]
-DiT-L/2 Random 100k 24.8 11.9 74.3 0.65 0.42
-DiT-L/2 D2C (Ours) 100k 14.8 6.9 109.2 0.63 0.52
-DiT-L/2 Random 300k 17.1 12.8 130.6 0.64 0.41
-DiT-L/2 D2C (Ours) 300k 5.8 15.1 318.9 0.77 0.29
-SiT-L/2 Random 100k 13.3 22.8 197.1 0.69 0.68
-SiT-L/2 D2C (Ours) 100k 9.1 14.3 261.7 0.72 0.34
-SiT-L/2 Random 300k 5.0 13.6 316.9 0.76 0.27
-SiT-L/2 D2C (Ours) 300k 4.22 11.6 289.7 0.79 0.24
-Figure 11: Generated samples on ImageNet 512$\times$512 from SiT-L/2 trained with D2C using a 10K dataset (CFG=1.5).
-Appendix K Visualization
-Figure 12: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "macaw"(88)
-Figure 13: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "arctic wolf"(270)
-Figure 14: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "jaguar"(290)
-Figure 15: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "otter"(360)
-Figure 16: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "lesser panda"(387)
-Figure 17: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "panda"(388)
-Figure 18: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "fire truck"(555)
-Figure 19: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "cheeseburger"(933)
-Figure 20: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "lake shore"(975)
-Figure 21: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "volcano"(980)
+Visual samples in Fig.
+11 further confirm the high fidelity and diversity of generations at $\times512$ resolution, demonstrating that D2C generalizes effectively to high-resolution settings.
+Appendix M Visualization
+Figure 12: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”macaw”(88)
+Figure 13: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”arctic wolf”(270)
+Figure 14: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”jaguar”(290)
+Figure 15: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”otter”(360)
+Figure 16: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”lesser panda”(387)
+Figure 17: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”panda”(388)
+Figure 18: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”fire truck”(555)
+Figure 19: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”cheeseburger”(933)
+Figure 20: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”lake shore”(975)
+Figure 21: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”volcano”(980)